### v1

# QE4PE: Word-level Quality Estimation for Human Post-Editing

Abstract
Word-level quality estimation (QE) detects erroneous spans in machine translations, which can direct and facilitate human post-editing. While the accuracy of word-level QE systems has been assessed extensively, their usability and downstream influence on the speed, quality and editing choices of human post-editing remain understudied. Our QE4PE study investigates the impact of word-level QE on machine translation (MT) post-editing in a realistic setting involving 42 professional post-editors across two translation directions. We compare four error-span highlight modalities, including supervised and uncertainty-based word-level QE methods, for identifying potential errors in the outputs of a state-of-the-art neural MT model. Post-editing effort and productivity are estimated by behavioral logs, while quality improvements are assessed by word- and segment-level human annotation. We find that domain, language and editors’ speed are critical factors in determining highlights’ effectiveness, with modest differences between human-made and automated QE highlights underlining a gap between accuracy and usability in professional workflows.
1 Introduction
Recent years saw a steady increase in the quality of machine translation (MT) systems and their widespread adoption in professional translation workflows (Kocmi et al., 2024a). Still, human post-editing of MT outputs remains a fundamental step to ensure high-quality translations, particularly for challenging textual domains requiring native fluency and specialized terminology (Liu et al., 2024). Quality estimation (QE) techniques were introduced to reduce post-editing effort by automatically identifying problematic MT outputs without the need for human-written reference translations and were quickly integrated into industry platforms (Tamchyna, 2021). Segment-level QE models correlate well with human perception of quality (Freitag et al., 2024) and exceed the performance of reference-based metrics in specific settings (Rei et al., 2021; Amrhein et al., 2022, 2023). On the other hand, word-level QE methods for identifying error spans requiring revision have received less attention in the past due to their modest agreement with human annotations, despite their promise for more granular and interpretable quality assessment in line with modern MT practices (Zerva et al., 2024). In particular, while the accuracy of these approaches is regularly assessed in evaluation campaigns, research has rarely focused on assessing the impact of such techniques in realistic post-editing workflows, with notable exceptions suggesting limited benefits (Shenoy et al., 2021; Eo et al., 2022). This hinders current QE evaluation practices: by foregoing experimental evaluation with human editors, it is implicitly assumed that word-level QE will become helpful once sufficient accuracy is reached, without accounting for the additional challenges towards a successful integration of these methods in post-editing workflows.
In this study dubbed QE4PE (Quality Estimation for Post Editing), we address this gap by conducting a large-scale study with 42 professional translators for the EnglishItalian and EnglishDutch directions to measure the impact of word-level QE on editing quality, productivity and usability. We aim for a realistic and reproducible setup, employing the high-quality open-source NLLB 3.3B MT model (NLLB Team et al., 2024) to translate challenging documents from biomedical and social media domains. We then conduct a controlled evaluation of post-editing with error spans using four highlight modalities: a supervised state-of-the-art QE model trained on human error annotations (XCOMET, Guerreiro et al., 2024a), an unsupervised method leveraging the uncertainty of the MT model during generation, oracle error spans obtained from the consensus of previous human post-editors, and a no highlight baseline. The human post-editing is done in a simple online interface built for this study, GroTE, supporting the real-time logging of granular editing data for a quantitative assessment of editing effort and productivity across highlight modalities. We also survey professionals using an online questionnaire to collect qualitative feedback about the usability and quality of the MT model, as well as the interface and error span highlights. Finally, a subset of the original MT outputs and their post-edited variants is annotated following the MQM and ESA protocols (Lommel et al., 2014; Kocmi et al., 2024b) to verify quality improvements after post-editing. See Figure 1 for an overview of the study. More generally, our work embraces recent calls for an evaluation of translation technologies that is centered on users’ experience (Guerberof-Arenas and Moorkens, 2023; Savoldi et al., 2025).
We release all data, code and the GroTE editing interface to foster future studies on the usability of error span highlighting techniques for other word-level QE methods and translation directions.111https://github.com/gsarti/qe4pe
2 Related Work
MT Post-editing
Human post-editing of MT outputs is increasingly common in professional translator workflows, as it was shown to increase the productivity of translators while preserving translation quality across multiple domains (Liu et al., 2024). However, many factors were found to influence the variability of post-editing productivity across setups, including MT quality (Zouhar et al., 2021b), interface familiarity (Läubli et al., 2021), individual variability and source-target languages typological similarity (Sarti et al., 2022). Studies evaluating the post-editing process generally focus on productivity, i.e. number of processed words/characters per minute, and the temporal, technical and cognitive dimensions of post-editing effort, operationalized through behavioral metrics such as editing time, keystrokes and pauses (Krings, 2001; Sarti et al., 2022). We adopt these metrics for the QE4PE study and relate them to different highlight modalities.
Quality estimation for MT
The field of quality estimation was initially concerned with MT model uncertainty (Blatz et al., 2004; Specia et al., 2009), but in time began focusing on predicting translation quality even without using references (Turchi et al., 2013, 2014; Kepler et al., 2019; Thompson and Post, 2020 inter alia). Advances in segment- and word-level QE research are regularly assessed in annual WMT campaigns (Fomicheva et al., 2021; Zerva et al., 2022, 2024; Blain et al., 2023), where the best-performing QE systems are usually Transformer-based language models trained on human quality judgments, such as the popular COMET model suite (Rei et al., 2020, 2021, 2022). The widespread adoption of the fine-grained Multidimensional Quality Metrics scale (MQM, Lommel et al., 2014) prompted a paradigm shift in MT evaluation, leading to new QE metrics predicting quality at various granularity levels (Kocmi and Federmann, 2023; Fernandes et al., 2023; Guerreiro et al., 2024a). Aside from supervised models, unsupervised methods exploiting model uncertainty and its internal mechanisms were proposed as efficient alternatives to identify potential error spans for MT (Fomicheva et al., 2020; Dale et al., 2023; Xu et al., 2023; Himmi et al., 2024, surveyed by Leiter et al. 2024) and other language generation tasks (Kuhn et al., 2023; Farquhar et al., 2024). In this work, we compare the downstream effectiveness of state-of-the-art supervised and unsupervised word-level QE metrics for post-editing settings.
Human Assessment of Quality Estimation
Automatic QE methods are widely used in the translation industry for triaging automatic translations (Tamchyna, 2021). While QE usage has been found helpful to increase the confidence and speed of human assessment (Mehandru et al., 2023; Zouhar et al., 2024b), an incautious usage of these techniques can lead to users being misled (Zouhar et al., 2021a). Interfaces supporting word-level error highlights were developed for studying MT post-editing (Coppers et al., 2018; Herbig et al., 2020) and code reviewing (Sun et al., 2022; Vasconcelos et al., 2024), with results suggesting that striking the right balance of user-provided information is fundamental to improve the editing experience and prevent cognitive overload. Most similar to our study, Shenoy et al. (2021) investigated the effect of synthetic word-level QE highlights for EnglishGerman post-editing on Wikipedia data, concluding that current word-level QE accuracy was still insufficient to produce tangible productivity benefits. In this work, we expand the scope of such evaluation by including two translation directions, two challenging real-world text domains and state-of-the-art MT and QE systems and methods.
3 Experimental Setup
3.1 Structure of the Study
Our study is organized in five stages:
1) Oracle Post-editing
As a preliminary step, segments later used in the main assessment are post-edited by three professionals per direction using their preferred interface without logging. This allows us to obtain post-edits and produce oracle word-level spans based on the editing consensus of multiple human professionals. Translators involved in this stage are not involved further in the study.
2) Pretask (Pre)
The pretask allows the core translators (12 per language direction, see Section 3.4) to familiarize themselves with the GroTE interface and text highlights. Before starting, all translators complete a questionnaire to provide demographic and professional information about their profile (Table 7). In the pretask, all translators work in an identical setup, post-editing a small set of documents similar to those of the main task with supervised highlights. We rank the translators into three groups (faster, average, slower) based on their speed from editing logs. Individuals from each group are then assigned randomly to highlight modalities to ensure an equal representation of editing speeds in each highlight modality. This procedure is repeated independently for both translation directions to balance subjective variability in the main task.
3) Main Task (Main)
This task, conducted in the two weeks following the pretask, covers the majority of the collected data and is the main object of study for the analyses of Section 4. In the main task, 24 core translators work on the same texts using the GroTE interface, but each with their assigned highlight modality (three translators per modality in each translation direction, as in Figure 1). After the main task, translators complete a questionnaire on the quality and usability of the MT outputs, the interface and, where applicable, word highlights.
4) Post-Task (Post)
After Main, core translators are asked to post-edit an additional small set of related documents with GroTE, but this time working all with the no highlight modality. This step lets us obtain baseline editing patterns for each translator to estimate individual speed and editing differences across highlight modalities without the confounder of interface proficiency of Pre.
5) Quality Assessment (QA)
Finally, a subset of 148 main task segments is selected randomly for manual annotation by six additional translators per direction (see Section 3.4). For each segment, the original MT output and all its post-edited versions are annotated with MQM error spans, including minor/major error severity and MQM error categories (Lommel et al., 2014). Moreover, the annotator proposes corrections for each error span, ultimately providing a 0-100 quality score. We adopt this scoring system, which closely follows the ESA protocol (Kocmi et al., 2024b), following recent results showing its effectiveness and efficiency.222See Figure 6 for an overview of setup and guidelines.
In summary, for each translation direction we collect 3 full sets of oracle post-edits, 12 full sets of edits with behavioral logs for pre, main and post task data, 12 pre- and post-task questionnaire responses, and 13 subsets of main task post-edits/MT outputs annotated with MQM error spans, corrections and segment-level ESA ratings.
3.2 Highlight Modalities
We conduct our study on four highlight modalities across two severity levels (minor and major errors). Using multiple severity levels follows the current MT evaluation practices (Freitag et al., 2021, 2024), and previous results showing that users tend to prefer more informative word-level highlights (Shenoy et al., 2021; Vasconcelos et al., 2024).
No Highlight
The text is presented as-is, without any highlighted spans. This setting serves as a baseline to estimate the default post-editing quality and productivity using our interface.
Oracle
Following the Oracle Post-editing phase, we produce oracle error spans from the editing consensus of human post-editors. We label text spans that were edited by two out of three translators as minor, and those edited by all three translators as major, following the intuition that more critical errors are more likely to be identified by several annotators, while minor changes will show more variance across subjects. This modality serves as a best-case scenario, providing an upper bound for future improvements in word-level QE quality.
Supervised
In this setting, word-level error spans are obtained using XCOMET-XXL (Guerreiro et al., 2024a), which is a multilingual Transformer encoder (Goyal et al., 2021) further trained for joint word- and sentence-level QE prediction. We select XCOMET-XXL in light of its broad adoption, open accessibility and state-of-the-art performance in QE across several translation directions (Zerva et al., 2024). For the severity levels, we use the default labels predicted by the model, ignoring confidence scores and mapping critical labels to the major level.
Unsupervised
In this modality, we exploit the access to the MT model producing the original translations to obtain uncertainty-based highlights. We assess the agreement of two unsupervised QE methods employing the MT model log-probabilities with human post-edits, following recent results showing these can act as significant predictors of human editing time, and hence of the presence of potential issues (Lim et al., 2024). Scores were collected for the EnglishItalian and EnglishDutch directions of QE4PE Oracle post-edits and DivEMT (Sarti et al., 2022) to identify the best-performing method.333Unsupervised evaluation results are shown in Table 13. Token-level highlights are aggregated at the word level to match the granularity of other modalities. Results show a strong performance for the variance of next-token log-probabilities for 10 steps of Monte Carlo Dropout (Gal and Ghahramani, 2016), even surpassing the agreement of supervised metrics across both datasets. We ultimately select this method, which was previously shown to provide an efficient alternative to model ensembling for estimating epistemic uncertainty at the sentence level (Fomicheva et al., 2020).
3.3 Data and MT model
MT Model
On the one hand, the MT model must achieve high translation quality in the selected languages to ensure our experimental setup applies to state-of-the-art proprietary systems. Still, the MT model should be open-source and have a manageable size to ensure reproducible findings and enable the computation of uncertainty for the unsupervised setting. All considered, we use NLLB 3.3B (NLLB Team et al., 2024), a widely-used MT model achieving industry-level performances across 200 languages (Moslem et al., 2023).
Data selection
We begin by selecting two translation directions, EnglishItalian and EnglishDutch. We intentionally focus on out-of-English translations as they are generally more challenging for modern MT models (Kocmi et al., 2023). We aim to identify documents that are manageable for professional translators without domain-specific expertise but still prove challenging for our MT model to ensure a sufficient amount of error spans across modalities.
We begin by translating 3672 multi-segment English documents from the WMT23 General and Biomedical MT shared tasks (Kocmi et al., 2023; Neves et al., 2023) and MT test suites to Dutch and Italian. Then, XCOMET-XXL is used to produce a first set of segment-level QE scores and word-level error spans for all segments. To make the study tractable, we further narrow down the selection of documents according to several heuristics to ensure a realistic editing experience and a balanced occurrence of error spans (see Appendix A for details and data statistics). This procedure yields 351 documents, from which we manually select a subset of 64 documents (413 segments, 8744 source words per post-editor) across two domains:
Biomedical abstracts extracted from PubMed from the WMT23 Biomedical Translation Task (Neves et al., 2023), including domain-specific terminology.
While the presence of multiple domains in the same task can render our post-editing setup less realistic, we deem it essential to test the cross-domain validity of our findings.
Critical Errors
Before producing highlights, we manually introduce 13 critical errors in main task segments to assess post-editing thoroughness. Errors are produced, for example, by negating statements, inverting the polarity of adjectives, inverting numbers, and corrupting acronyms. We replicate the errors in both translation directions to enable direct comparison. Most of these errors were correctly identified across all three highlight modalities (examples in Table 4).
3.4 Participants
For both directions, professional translation providers Translated and Global Textware recruited three translators for the Oracle post-editing stage, 12 for Pre, Main and Post tasks, and finally six for the QA stage. All translators were freelancers with native proficiency in their target language and self-assessed proficiency of at least C1 in English. Almost all translators had more than two years of professional translation experience and regularly post-edited MT outputs (details in Table 7).
3.5 Editing Interface
We develop a custom interface, which we name Groningen Translation Environment (GroTE, Figure 2), to support editing over texts with word-level highlights. While the MMPE tool used by Shenoy et al. (2021) provide extensive multimodal functionalities (Herbig et al., 2020), we aim for a bare-bones setup to avoid confounders in the evaluation. GroTE is a web interface based on Gradio (Abid et al., 2019) and hosted on the Hugging Face Spaces to enable multi-user data collection online. Upon loading a document, source texts and MT outputs for all segments are presented in two columns following standard industry practices. For modalities with highlights, the interface provides an informative message and supports the removal of all highlights on a segment via a button, with highlights on words disappearing automatically upon editing, as in Shenoy et al. (2021). The interface supports real-time logging of user actions to analyze the editing process. In particular, we log the start and end times for each edited document, the accessing and exiting of segment textboxes, highlight removals, and individual keystrokes during editing.
GroTE intentionally lacks standard features such as translation memories, glossaries, and spellchecking to ensure equal familiarity among translators, ultimately controlling for editor proficiency with these tools, as done in previous studies (Shenoy et al., 2021; Sarti et al., 2022). While most translators noted the lack of features in our usability assessment, the majority also found the interface easy to set up, access, and use (Table 7).
4 Analysis
4.1 Productivity
We obtain segment- and document-level edit times and compute editing productivity as the number of processed source characters per second over the sum of all document-level edit times. To account for potential breaks taken by post-editors during editing, we filter out pauses between logged actions longer than 5 minutes. We note that our findings are not sensitive to this filtering.
Do Highlights Make Post-editors Faster?
Figure 3 shows translators’ productivity across stages, with every dot corresponding to the productivity of a single individual. We observe that no highlight modality leads to systematically faster editing across all speed groups and that the ordering of Pre-task speed groups is maintained in the following stages despite the different highlight modalities. These results suggest that individual variability in editing speed is more critical than highlight modality in predicting editing speed. However, faster EnglishDutch translators achieve outstanding productivity ( char/min) only for No Highlight, and, Oracle modalities, suggesting that lower-quality highlights hinder editing speed.
We validate these observations by fitting a negative binomial mixed-effect model on segment-level editing times (model details in Table 5). Excluding random factors such as translator and segment identity from the model produces a significant drop in explained variance, confirming the inherent variability of editing times (). Model coefficients show that MT output length and the proportion of highlighted characters are the main factors driving an increase in editing times, possibly reflecting an increase in cognitive effort to process additional information. We find highlights to have a significant impact on increasing the editing speed of EnglishItalian translators , but a minimal impact for EnglishDutch. Comparing the productivity of the same translator editing with and without highlights (Main vs Post), two-thirds of the translators editing with highlights were up to two times slower on biomedical texts. However, the same proportion of translators was up to three times faster on social media texts across both directions.
In summary, we find that highlight modalities are not predictive of edit times on their own, but translation direction and domain play an important role in determining the effect of highlights on editing productivity. We attribute these results to two main factors, which will remain central in the analysis of the following sections: (1) the different propensity of translators to act upon highlighted issues in the two tested directions, and (2) the different nature of errors highlighted across domains.
4.2 Highlights and Edits
We then examine how highlights are distributed across modalities and how they influence the editing choices of human post-editors.
| Base Freq. | Measured | Projected | ||||||||
| EnglishItalian | ||||||||||
| No High. | - | 0.05 | - | - | - | - | - | - | - | - |
| Random | 0.16 | - | - | - | - | - | 0.06 | 1.20 | 0.18 | 1.20 |
| Oracle | 0.15 | 0.12 | 0.37 | 4.62 | 0.45 | 4.1 | 0.18↓0.19 | 6.00↑1.38 | 0.55↑0.10 | 4.23↑0.14 |
| Unsup. | 0.16 | 0.13 | 0.25 | 2.27 | 0.21 | 2.2 | 0.11↓0.14 | 2.75↑0.48 | 0.37↑0.16 | 2.47↑0.26 |
| Sup. | 0.12 | 0.16 | 0.28 | 2.00 | 0.22 | 2.0 | 0.14↓0.14 | 3.50↑1.50 | 0.35↑0.13 | 3.18↑1.18 |
| EnglishDutch | ||||||||||
| No High. | - | 0.14 | - | - | - | - | - | - | - | - |
| Random | 0.17 | - | - | - | - | - | 0.16 | 1.14 | 0.19 | 1.19 |
| Oracle | 0.20 | 0.10 | 0.26 | 4.33 | 0.53 | 3.12 | 0.28↑0.02 | 2.55↓1.78 | 0.40↓0.13 | 2.35↓0.77 |
| Unsup. | 0.20 | 0.11 | 0.20 | 2.50 | 0.36 | 2.00 | 0.22↑0.02 | 1.83↓0.67 | 0.31↓0.05 | 1.72↓0.28 |
| Sup. | 0.12 | 0.09 | 0.24 | 3.43 | 0.33 | 3.30 | 0.28↑0.04 | 2.33↓1.10 | 0.24↓0.09 | 2.40↓0.90 |
Agreement Across Modalities
First, we quantify how different modalities agree in terms of highlights’ distribution and editing. We find that cross-modality highlight overlap ranges between 15% and 39%, with the highest overlap for EnglishItalian social media and EnglishDutch biomedical texts.444Scores are normalized to account for highlight frequencies across modalities. Highlight agreement shown in Table 8. Despite the relatively low highlight agreement, we find an average agreement of 73% for post-edited characters across modalities. This suggests that edits are generally uniform regardless of highlight modalities and are not necessarily restricted to highlighted spans.555Editing agreement is shown in Figure 8.
Do Highlights Accurately Identify Potential Issues?
Table 1 (Base Freq.) shows raw highlight and edit frequencies across modalities. We observe different trends across the two language pairs: for EnglishItalian, post-editors working with highlights edit more than twice as much as translators with No Highlight, regardless of the highlight modality. On the contrary, for EnglishDutch they edit 33% less in the same setting. These results suggest a different attitude towards acting upon highlighted potential issues across the two translation directions, with EnglishItalian translators appearing to be conditioned to edit more when highlights are present. We introduce four metrics to quantify highlights-edits overlap:
and , reflecting highlights’ precision and recall in predicting edits, respectively.
shows how much more likely an edit is to fall within rather than outside highlighted characters.
shows how much more likely it is for a highlight to mark edited rather than unmodified spans.
Intuitively, character-level recall should be more indicative of highlight quality compared to precision , provided that word-level highlights can be useful even when not minimal.666For example, if the gender of the highlighted word traduttore is changed to the feminine traduttrice, (edit correctly and fully predicted) but since word stem characters are left unchanged. Table 1 (Measured) shows metric values across the three highlight modalities (breakdowns by domain and speed shown in Tables 10 and 11). As expected, Oracle highlights obtain the best performance in terms of precision and recall, with , in particular, being significantly higher than the other two modalities across both directions.
Surprisingly, we find no significant precision and recall differences between Supervised and Unsupervised highlights despite the word-level QE training of XCOMET used in the former modality. Moreover, they support the potential of unsupervised, model internals-based techniques to complement or substitute more expensive supervised approaches. Still, likelihood ratios for all modalities and directions indicate that highlights are 2-4 times more likely to precisely and comprehensively encompass edits than non-highlighted texts. This suggests that even imperfect highlights that do not reach Oracle-level quality might effectively direct editing efforts toward potential issues. We validate these observations by fitting a zero-inflated negative binomial mixed-effects model to predict segment-level edit rates. Results confirm a significantly higher edit rate for EnglishItalian highlighted modalities and the social media domain with (features and significances shown in Appendix Table 6). We find a significant zero inflation associated with translator identity, suggesting the choice of leaving MT outputs unedited is highly subjective.
Do Highlights Influence Editing Choices?
Having observed a clear influence of highlights on the editing rate of translators, we question whether the relatively high and values for some highlight modalities are not artificially inflated by the eagerness of translators to intervene on highlighted spans, rather than reflecting actual highlight quality. In other words, do highlights identify actual issues, or do they condition translators to edit when they otherwise would not? To answer this, we propose to project highlights from a selected modality—in which highlights were shown during editing—onto the edits performed by the No Highlight translators on the same segments. The resulting difference between measured and projected metrics can then be taken as an estimate for the impact of highlight presentation on their resulting accuracy. To further ensure the soundness of our analysis, we use a set of projected Random highlights as a lower bound for highlight performance. To make the comparison fair, Random highlights are created by randomly highlighting words in MT outputs matching the average word-level highlight frequency across all highlighted modalities given the current domain and translation direction. Table 1 (Projected) shows results for the three highlighted modalities. Projected precision scores are influenced by edit frequency, and hence see a major decrease for EnglishItalian, where the No Highlight edit rate is much lower. However, the increase in across all EnglishItalian modalities confirms that, despite the lower edit proportion, highlighted texts remain notably more likely to be edited than non-highlighted ones. Conversely, the lower for EnglishDutch shows that edits become much less skewed towards highlighted spans in this direction when accounting for presentation bias. and , which are not influenced by different edit rates, exhibit the same trend. Finally, projected metrics remain consistently above the Random baseline, suggesting a higher-than-chance ability to identify errors even in the worst-performing settings.
Overall, while the presence of highlights makes EnglishItalian translators more likely to intervene in MT outputs, their location in the MT output often pinpoints issues that would be edited regardless of highlighting. EnglishDutch translators, on the contrary, intervene at roughly the same rate regardless of highlights presence, but their edits are focused mainly on highlighted spans when they are present. This difference is consistent across all subjects in the two directions despite the identical setup and comparable MT and QE quality across languages. This suggests that cultural factors might play a non-trivial role in determining the usability and influence of QE methods regardless of span accuracy, a phenomenon previously observed in human-AI interaction studies (Ge et al., 2024).
4.3 Quality Assessment
We continue our assessment by inspecting the quality of MT and post-edited outputs along three dimensions. First, we use XCOMET segment-level QE ratings as an automatic approximation of quality and compare them to human-annotated quality scores collected in the last phase of our study. For efficiency, these are obtained for the 0-100 Direct Assessment scale commonly used in QE evaluation (Specia et al., 2020), but following an initial step of MQM error annotation to condition scoring on found errors, as prescribed by the ESA protocol (Kocmi et al., 2024b). Secondly, MQM error span annotations are used to analyze the distribution of error categories. Thirdly, we manually assess critical errors, which we inserted to quantify whether highlight modalities can help address unambiguous issues.
Do Highlights Influence Post-Editing Quality?
In this stage, we focus particularly on edited quality improvements, i.e. the quality difference between MT and post-edited versions of the same segment. Positive scores for this metric reflect an increase in quality after the post-editing stage and are bounded by the maximal achievable quality gain given the initial MT quality. Figure 4 shows median improvement values across quality bins defined from the distribution of initial MT quality scores (shown in histograms), in which all post-edited versions of each MT output appear as separate observations. Positive median scores confirm that post-edits generally lead to quality improvements across all tested settings. However, we observe different trends across the two metrics: across both domains, XCOMET underestimates the human-assessed ESA quality improvement, especially for biomedical texts where it shows no improvements regardless of the initial MT quality. These results echo recent findings cautioning users against the poor performance of trained MT metrics for unseen domains and high-quality translations (Agrawal et al., 2024; Zouhar et al., 2024a). Focusing on the more reliable ESA scores, we observe large quality improvements from post-editing, as shown by near-maximal quality gains across most bins and highlight modalities. While No Highlight seems to underperform other modalities in the social media domain, the lack of more notable differences suggests that highlights’ quality impact might not be evident in terms of segment-level quality, motivating our next steps in the quality analysis.
We also find no clear relationship between translator speed and edited quality improvements, suggesting that higher productivity does not come at a cost for faster translators (Figure 10). This finding further confirms that neglecting errors is not the cause of the differences in editing patterns shown in previous sections.
Which Error Types Do Highlights Identify?
Figure 5 shows a breakdown of MQM annotations across modalities, domains and directions for three error macro-categories.777Detailed MQM error counts are shown in Table 9. At this granularity, differences across modalities become visible, with overall error counts showing a clear relation to from Table 1 (Oracle being remarkably better for EnglishItalian, with milder and more uniform trends in EnglishDutch). At least for EnglishItalian, these results confirm that an observable quality improvement from editing with highlights is present in the best-case Oracle scenario. We also note a different distribution of Accuracy and Style errors, with the former being more common in biomedical texts while the latter appearing more often for translated social media posts. We posit that differences in error types across domains might be the culprit behind the opposite productivity trends observed in Section 4.1: highlighted accuracy errors might lead to time-consuming terminology checks in biomedical texts. In contrast, style errors with simple fixes might be corrected more quickly in social media texts.
Do Highlights Detect Critical Errors?
We examine whether the critical errors we inserted were detected by different modalities, finding that while most modalities fare decently with more than 62% of critical errors highlighted, Unsupervised is the only setting for which all errors are correctly highlighted across both directions. Then, critical errors are manually verified in all outputs, finding that 16-20% more critical errors are edited in highlighted modalities compared to No Highlight (full results in Table 9). Hence, the presence of highlights might result in narrow but tangible quality improvements that remain undetected in coarser assessments, and finer-grained evaluations might be needed to quantify future improvements in word-level QE quality.
4.4 Usability
In post-task questionnaire answers (Table 2), most translators stated that MT outputs had average-to-high quality and that provided texts were challenging to translate. Highlights were generally found decently accurate, but they were generally not found useful to improve either productivity or quality (including Oracle ones). Interestingly, despite the convincing gains for critical errors measured in the last section, most translators stated that highlights did not influence their editing and did not help them identify errors that would have otherwise been missed. Concretely, this suggests that the potential quality improvements might not be easily perceived by translators and might have secondary importance compared to the extra cognitive load elicited by highlighted spans. When asked to comment about highlights, several translators called them “more of an eye distraction, as they often weren’t actual mistakes” and “not quite accurate enough to rely on them as a suggestion”. Some translators also stated that missed errors led them to “disregarding the highlights to focus on checking each sentence”. Despite their high quality, only one editor working with Oracle highlights found highlights helpful in “making the editing process faster and somehow easier”. Taken together, these comments convincingly point to a negative perception of the quality and usefulness of highlights, suggesting that improvement in QE accuracy may not be sufficient to improve QE usefulness in editors’ eyes.
| Question | EngIta | EngNld |
| MT outputs were generally of high quality. | ||
| Provided texts were challenging to translate. | ||
| Highlights … | ||
| … were generally accurate in detecting potential issues. | ||
| … were generally useful during editing. | ||
| … improved my editing productivity. | ||
| … improved the quality of my translations. | ||
| … required additional editing effort on my part. | ||
| … influenced my editing choices. | ||
| … helped identify errors I’d have otherwise missed. |
5 Conclusion
This study evaluated the impact of various error-span highlighting modalities, including automatic and human-made ones, on the productivity and quality of human post-editing in a realistic professional setting. Our findings highlight the importance of domain, language and editors’ speed in determining highlights’ effect on productivity and quality, underscoring the need for broad evaluations encompassing diverse settings. The limited gains of human-made highlights over automatic QE and their indistinguishable perception from editors’ assessment indicate that further gains in the accuracy of these techniques might not be the determining factor in improving their integration into post-editing workflows. In particular, future work might explore other directions to further assess and improve the usability of word-level QE highlights, for example, studying their impact on non-professional translators and language learners or combining them with edit suggestions to justify the presence of error spans.
Acknowledgements
Gabriele Sarti and Arianna Bisazza acknowledge the support of Imminent, which funded the data collection for this study with a Language Technology Grant. The authors are grateful to Translated S.r.l. and Global Textware B.v. and all translators that collaborated in the data collection process. Gabriele Sarti, Grzegorz Chrupała and Arianna Bisazza also acknowledge the support of the Dutch Research Council (NWO) for the project InDeep (NWA.1292.19.399). Arianna Bisazza is further supported by the NWO Talent Programme (VI.Vidi.221C.009). Ana Guerberof Arenas is supported by the European Research Council (ERC CoG INCREC 101086819).
Broader Impact and Ethical Considerations
Our study explicitly centers the experience of professional translators, responding to recent calls for user-centered evaluation of translation technologies. By prioritizing translators’ perspectives and productivity, we aim to contribute to methods that complement rather than replace human expertise. Our findings highlight a gap between user perception and measured quality improvements, suggesting that future efforts should focus primarily on improving the usability of these methods in editing interfaces. In particular, new assistive approaches for post-editing should not only strive to increase productivity, but rather reduce the cognitive burden associated with post-editing work. This insight is crucial for designing more user-centered quality estimation tools that genuinely support human work. Finally, our results underscores the importance of culturally adaptive design for assessing the effectiveness of proposed methodologies across participants with different cultural backgrounds. All participants in this study were professional translators who provided informed consent. The research protocol ensured anonymity and voluntary participation, with translators recruited through professional translation providers. The study’s open data release further promotes transparency, allowing other researchers to reproduce and extend our findings.
Appendix A Filtering Details and Statistics for QE4PE Data
1.
Documents should contain between 4 and 10 segments, each containing 10-100 words (959 docs). This ensures that all documents are roughly uniform in terms of size and complexity to maintain a steady editing flow (Section 3.5).
2.
The average segment-level QE score predicted by XCOMET-XXL is between 0.3 and 0.95, with no segment below 0.3 (429 docs). This forces segments to have a decent but still imperfect quality, excluding fully wrong translations.
3.
At least 3 and at most 20 errors spans per document, with no more than 30% of words in the document being highlighted (351 docs). This avoids overwhelming the editor with excessive highlighting, while still ensuring error presence.
The same heuristics were applied to both translation directions, selecting only documents matching our criteria in both cases. Table 3 provide some statistics for the selected examples.
| Task | Domain | # Docs | # Seg. | # Words |
| Pre | Social | 4 | 23 | 539 |
| Biomed. | 2 | 15 | 348 | |
| Main | Social | 30 | 160 | 3375 |
| Biomed. | 21 | 165 | 3384 | |
| Post | Social | 6 | 34 | 841 |
| Biomed. | 2 | 16 | 257 | |
| Total | 64 | 413 | 8744 | |
| Remove negation (13-6) | |
| English | No significant differences were found with respect to principal diagnoses […] |
| Dutch | Er werden geen significante verschillen significante verschillen gevonden met betrekking tot de belangrijkste diagnoses […] |
| Title literal translation (16-3) | |
| English | The Last of Us is an easy and canonical example of dad-ification. […] |
| Italian | The Last of Us L’ultimo di noi è un esempio facile e canonico di dad-ification. […] |
| Wrong term (48-5) | |
| English | […], , except for alkaline phosphatase. |
| Italian | […], ad eccezione della fosfatasi alcalina chinasi proteica. |
| Target: Seg. Edit Time, 5s bins from 0 to 600s | ||
| Feature | Coeff. | Significance |
| (Intercept) | 1.67 | *** |
| MT Num. Chars | 2.42 | *** |
| Highlight Ratio % | 1.59 | *** |
| Target Lang.: ITA | -0.34 | *** |
| Text Domain: Social | 0.31 | *** |
| Oracle Highlight | -0.79 | . |
| Sup. Highlight | 0.02 | |
| Unsup. Highlight | -0.07 | |
| MT XCOMET QE Score | 0.01 | *** |
| ITA:Oracle | 0.91 | *** |
| ITA:Sup. | 1.18 | *** |
| ITA:Unsup. | 0.48 | *** |
| Social:Oracle | -0.19 | ** |
| Social:Sup. | -0.34 | *** |
| Social:Unsup. | -0.22 | *** |
| Highlight Ratio:Oracle | -0.83 | * |
| Highlight Ratio:Sup. | -1.33 | *** |
| Edit Order | Random Factors | |
| Translator ID | ||
| Segment ID | ||
| Target: % of edited characters in a segment (0-100). | ||
| Feature | Coeff. | Significance |
| (Intercept) | 21.0 | *** |
| MT Num. Chars | 10.3 | *** |
| Highlight Ratio % | 7.1 | *** |
| Target Lang.: ITA | -9.9 | *** |
| Text Domain: Social | 10.9 | *** |
| Oracle Highlight | -5.2 | |
| Sup. Highlight | -4.7 | |
| Unsup. Highlight | -0.9 | |
| ITA:Oracle | 12.2 | *** |
| ITA:Sup. | 15.9 | *** |
| ITA:Unsup. | 13.4 | *** |
| Social:Oracle | 3.5 | *** |
| Social:Sup. | -0.4 | |
| Social:Unsup. | 2.1 | ** |
| Highlight Ratio:Oracle | -0.18 | |
| Highlight Ratio:Sup. | -1.78 | *** |
| Edit Order | Random Factors | |
| Translator ID | ||
| Segment ID | ||
| MT Num. Chars | Zero-Inflation Factors | |
| Target Lang | ||
| Text Domain | ||
| Translator ID | ||
| Identifier | Job | Eng. Lvl | Trans. YoE | Post-edit YoE | Post-edit % | Adv. CAT YoE | MT good/bad for: | Post-edit comment |
| eng-ita-nohigh-fast | Freelance (FT) | C1 | 2-5 | 2-5 | 100% | Often | Good: Productivity, quality, repetitive work. | PE better than from scratch when consistency is needed. |
| eng-ita-nohigh-avg | Freelance (PT) | C1 | >10 | <2 | 20% | Often | Good: Productivity, repetitive work. Bad: less creative. | PE produces unnatural sentences. |
| eng-ita-nohigh-slow | Freelance (PT) | C2 | >10 | 2-5 | 40% | Sometimes | Good: creativity. | Good for time saving. |
| eng-ita-oracle-fast | Freelance (FT) | C2 | 5-10 | 2-5 | 60% | Sometimes | Good: Productivity, repetitive work. Bad: less creative. | Good for productivity, humans always needed. |
| eng-ita-oracle-avg | Freelance (FT) | C2 | 5-10 | 5-10 | 20% | Always | Good: productivity, terminology. | Good for tech docs, not for articulated texts. |
| eng-ita-oracle-slow | Freelance (FT) | C2 | 2-5 | 5-10 | 80% | Always | Good: Productivity, repetitive work. | Useful for consistency and productivity, unless creativity is needed. |
| eng-ita-unsup-fast | Freelance (FT) | C1 | <2 | <2 | 60% | Often | Good: Productivity, terminology. Bad: less creative. | Humans will always be needed in translation. |
| eng-ita-unsup-avg | Freelance (FT) | C1 | >10 | 2-5 | 60% | Often | Good: Productivity, repetitive work. Bad: less creative. | An opportunity for translators. |
| eng-ita-unsup-slow | Freelance (FT) | C1 | 5-10 | 5-10 | 80% | Always | Good: Productivity, repetitive work. Bad: less creative. | Good for focusing on detailed/cultural/creative aspects of translations. |
| eng-ita-sup-fast | Freelance (PT) | C1 | >10 | 2-5 | 40% | Often | Good: Productivity, quality, repetitive work, terminology. | Improves quality and consistency. |
| eng-ita-sup-avg | Freelance (FT) | C1 | >10 | 5-10 | 100% | Always | Good: Productivity, repetitive work. Bad: less creative. | Consistency improved, but less variance means less creativity. |
| eng-ita-sup-slow | Freelance (FT) | C1 | >10 | 2-5 | 20% | Always | Good: Productivity, creativity, quality, repetitive work. | Good for productivity, but does not work on creative texts. |
| eng-nld-nohigh-fast | Freelance (FT) | C1 | >10 | >10 | 40% | Often | Good: Productivity, terminology. Bad: creativity. | Widespread but still too literal |
| eng-nld-nohigh-avg | Freelance (FT) | C2 | >10 | 2-5 | 40% | Always | Good: Repetitive work. Bad: creativity, often wrong, worse quality. | Increase in productivity to save on costs brings down quality. |
| eng-nld-nohigh-slow | Freelance (FT) | C2 | >10 | 5-10 | 100% | Often | Good: Creativity, quality, repetitive work, terminology. | Working with MT can be creative beyond PE. |
| eng-nld-oracle-fast | Freelance (FT) | C1 | 5-10 | 5-10 | 80% | Always | Good: Productivity, quality, repetitive work, terminology. | Good for tech docs and repetition. |
| eng-nld-oracle-avg | Freelance (FT) | C2 | >10 | 2-5 | 40% | Always | Bad: less creative, less productive, often wrong | Bad MT is worse than no MT for specialized domains. |
| eng-nld-oracle-slow | Freelance (FT) | C2 | >10 | 2-5 | 60% | Often | Good: Productivity, repetitive work. Bad: cultural references. | More productivity at the cost of idioms and cultural factors. |
| eng-nld-unsup-fast | Freelance (FT) | C2 | 5-10 | 2-5 | 40% | Often | Good: all. Bad: often wrong, worse quality. | PE makes you less in touch with the texts and often poorly paid. |
| eng-nld-unsup-avg | Freelance (FT) | C2 | 5-10 | 2-5 | 60% | Sometimes | Good: Productivity, quality, repetitive work, terminology. Bad: wrong. | Practical but less effective for longer passages. |
| eng-nld-unsup-slow | Freelance (FT) | C2 | >10 | 2-5 | 40% | Always | Good: repetitive work, productivity, terminology | Improves consistency and productivity if applied well. |
| eng-nld-sup-fast | Freelance (FT) | C2 | >10 | 5-10 | 60% | Often | Good: repetitive work, creativity, terminology | Useful, but worries about job loss |
| eng-nld-sup-avg | Freelance (FT) | C2 | >10 | 10 | 60% | Sometimes | Good: terminology, creativity | Useful for inspiration on better translations |
| eng-nld-sup-slow | Freelance (FT) | C1 | 5-10 | 5-10 | 80% | Always | Good: repetitive work, productivity | Better productivity at the cost of creativity. |
| Identifier | Freq. Issues | MT quality | MT fluency | MT accuracy | High. accurate | High. useful | Interface clear | Task difficult | Highlights statements | |||||
| Speed? | Quality? | Effort? | Influence | Spot errors | Enjoy? | |||||||||
| eng-ita-nohigh-fast | inflection,additions,omissions | 4 | 0.8 | 0.8 | - | - | 5 | 1 | - | - | - | - | - | - |
| eng-ita-nohigh-avg | multiple | 3 | 0.6 | 0.4 | - | - | 2 | 4 | - | - | - | - | - | - |
| eng-ita-nohigh-slow | terminology,omissions | 3 | 0.8 | 0.8 | - | - | 1 | 5 | - | - | - | - | - | - |
| eng-ita-oracle-fast | inflection,terminology | 5 | 0.4 | 0.8 | 4 | 4 | 4 | 5 | 5 | 2 | 1 | 1 | 1 | 4 |
| eng-ita-oracle-avg | syntax,terminology,omissions,no context | 3 | 0.4 | 0.6 | 2 | 1 | 2 | 3 | 1 | 1 | 4 | 1 | 1 | 1 |
| eng-ita-oracle-slow | syntax,no context | 3 | 0.6 | 0.6 | 2 | 2 | 2 | 5 | 1 | 1 | 1 | 1 | 4 | 1 |
| eng-ita-unsup-fast | omissions | 3 | 0.8 | 0.6 | 3 | 2 | 4 | 5 | 3 | 3 | 3 | 2 | 2 | 2 |
| eng-ita-unsup-avg | syntax,terminology,no context | 3 | 0.6 | 0.6 | 3 | 3 | 3 | 5 | 2 | 3 | 2 | 1 | 1 | 3 |
| eng-ita-unsup-slow | syntax,inflection,terminology,omissions | 3 | 0.4 | 0.6 | 2 | 2 | 3 | 4 | 2 | 2 | 3 | 3 | 4 | 4 |
| eng-ita-sup-fast | syntax,terminology,no context | 3 | 0.4 | 0.4 | 2 | 1 | 2 | 2 | 1 | 1 | 3 | 1 | 2 | 2 |
| eng-ita-sup-avg | syntax,terminology,no context | 3 | 0.4 | 0.4 | 2 | 2 | 3 | 5 | 3 | 2 | 4 | 3 | 3 | 4 |
| eng-ita-sup-slow | syntax,terminology,omissions,no context | 3 | 0.6 | 0.6 | 2 | 2 | 1 | 2 | 2 | 1 | 1 | 4 | 4 | 1 |
| eng-nld-nohigh-fast | syntax,terminology,omissions,no context | 3 | 0.2 | 0.4 | - | - | 4 | 4 | - | - | - | - | - | - |
| eng-nld-nohigh-avg | syntax,terminology,omissions,no context | 2 | 0.4 | 0.6 | - | - | 4 | 5 | - | - | - | - | - | - |
| eng-nld-nohigh-slow | terminology,omissions,no context | 2 | 0.2 | 0.4 | - | - | 3 | 5 | - | - | - | - | - | - |
| eng-nld-oracle-fast | syntax,inflection,terminology | 3 | 0.6 | 0.6 | 2 | 1 | 3 | 2 | 2 | 2 | 2 | 1 | 1 | 1 |
| eng-nld-oracle-avg | syntax | 3 | 0.8 | 0.6 | 4 | 3 | 3 | 4 | 3 | 3 | 3 | 3 | 2 | 3 |
| eng-nld-oracle-slow | syntax,terminology | 3 | 0.6 | 0.4 | 3 | 1 | 3 | 4 | 1 | 1 | 1 | 1 | 1 | 3 |
| eng-nld-unsup-fast | terminology,additions,omissions | 3 | 0.6 | 0.8 | 3 | 2 | 4 | 4 | 1 | 3 | 1 | 1 | 2 | 1 |
| eng-nld-unsup-avg | multiple | 3 | 0.6 | 0.6 | 4 | 3 | 2 | 4 | 3 | 3 | 4 | 3 | 2 | 3 |
| eng-nld-unsup-slow | syntax,terminology,omissions | 1 | 0.4 | 0.4 | 2 | 4 | 1 | 4 | 4 | 4 | 3 | 2 | 2 | 3 |
| eng-nld-sup-fast | terminology,omissions,no context | 3 | 0.6 | 0.4 | 2 | 2 | 3 | 5 | 1 | 1 | 5 | 3 | 1 | 1 |
| eng-nld-sup-avg | syntax,additions,no context | 3 | 0.4 | 0.6 | 2 | 2 | 2 | 4 | 1 | 1 | 1 | 1 | 2 | 3 |
| eng-nld-sup-slow | multiple | 5 | 0.8 | 1 | 4 | 3 | 2 | 5 | 3 | 3 | 2 | 2 | 2 | 4 |
| Modalities | EnglishItalian | EnglishDutch | Both | |||||||
| Bio | Social | Both | Bio | Social | Both | Bio | Social | Both | ||
| Oracle and | Sup. | 0.17 | 0.32 | 0.25 | 0.38 | 0.29 | 0.34 | 0.26 | 0.29 | 0.29 |
| Unsup. | 0.14 | 0.30 | 0.20 | 0.31 | 0.27 | 0.28 | 0.22 | 0.29 | 0.24 | |
| Supervised and | Oracle | 0.19 | 0.31 | 0.26 | 0.30 | 0.26 | 0.29 | 0.24 | 0.29 | 0.28 |
| Unsup. | 0.19 | 0.33 | 0.25 | 0.28 | 0.24 | 0.25 | 0.24 | 0.29 | 0.25 | |
| Unsupervised and | Oracle | 0.22 | 0.32 | 0.27 | 0.35 | 0.30 | 0.33 | 0.28 | 0.31 | 0.30 |
| Sup. | 0.22 | 0.37 | 0.30 | 0.39 | 0.27 | 0.33 | 0.30 | 0.31 | 0.32 | |
| # Doc.-Seg. | Error Type | Has Highlight | % Post-edited | |||||
| Oracle | Unsup. | Sup. | No High. | Oracle | Unsup. | Sup. | ||
| 1-8 | Wrong number | nld | Both | Both | 67 | 83 | 83 | 83 |
| 13-6 | Remove negation | ita | Both | Both | 50 | 33 | 33 | 50 |
| 16-3 | Title literal translation | Both | Both | Both | 83 | 100 | 100 | 100 |
| 20-1 | Wrong acronym | nld | Both | ita | 0 | 33 | 33 | 33 |
| 20-7 | Wrong acronym (1) | Neither | Both | Neither | 0 | 58 | 50 | 25 |
| 20-7 | Wrong acronym (2) | nld | Both | ita | 0 | 58 | 50 | 25 |
| 22-1 | Name literal translation | Both | Both | Both | 50 | 50 | 83 | 67 |
| 23-4 | Addition | nld | Both | Neither | 100 | 100 | 83 | 50 |
| 31-2 | Wrong acronym | nld | Both | Neither | 17 | 33 | 17 | 33 |
| 34-7 | Numbers swapped | nld | Both | nld | 17 | 50 | 33 | 67 |
| 37-4 | Verb polarity inverted | Both | Both | Both | 67 | 83 | 67 | 83 |
| 43-5 | Wrong name | Both | Both | Both | 50 | 83 | 67 | 83 |
| 48-5 | Wrong term | nld | Both | nld | 67 | 50 | 83 | 83 |
| Total | 65 | 100 | 62 | 44 | 63 | 60 | 60 | |
| Domain | Modality | |||||||||
| EnglishItalian | ||||||||||
| Biomed. | Random | .12 | - | - /.02 | - /.02 | - /1.0 | - /.11 | - /.13 | - /0.8 | - /.03 |
| No High. | - | .02 | - | - | - | - | - | - | - | |
| Oracle | .08 | .07 | .26/.08 | .05/.02 | 5.2/4.0 | .30/.26 | .06/.08 | 5.0/3.2 | .28/.12 | |
| Unsup. | .16 | .10 | .18/.06 | .08/.02 | 2.2/3.0 | .29/.36 | .14/.15 | 2.0/2.4 | .22/.10 | |
| Sup. | .11 | .12 | .18/.05 | .11/.02 | 1.6/2.5 | .16/.23 | .10/.10 | 1.6/2.3 | .17/.08 | |
| Social | Random | .20 | - | - /.09 | - /.09 | - /1.0 | - /.21 | - /.20 | - /1.0 | - /.13 |
| No High. | - | .09 | - | - | - | - | - | - | - | |
| Oracle | .25 | .20 | .42/.23 | .13/.04 | 3.2/5.7 | .52/.66 | .18/.21 | 2.8/3.1 | .46/.34 | |
| Unsup. | .17 | .18 | .35/.19 | .14/.07 | 2.5/2.7 | .33/.37 | .14/.15 | 2.3/2.4 | .34/.25 | |
| Sup. | .15 | .21 | .38/.23 | .18/.06 | 2.1/3.8 | .27/.39 | .11/.12 | 2.4/3.2 | .32/.29 | |
| EnglishDutch | ||||||||||
| Biomed. | Random | .17 | - | - /.12 | - /.10 | - /1.2 | - /.19 | - /.17 | - /1.1 | - /.15 |
| No High. | - | .10 | - | - | - | - | - | - | - | |
| Oracle | .21 | .08 | .21/.20 | .05/.08 | 4.2/2.5 | .52/.41 | .18/.18 | 2.8/2.2 | .30/.27 | |
| Unsup. | .23 | .09 | .17/.17 | .07/.08 | 2.4/2.1 | .43/.38 | .21/.21 | 2.0/1.8 | .24/.23 | |
| Sup. | .12 | .08 | .20/.21 | .06/.09 | 3.3/2.3 | .30/.25 | .11/.11 | 2.7/2.2 | .24/.23 | |
| Social | Random | .16 | - | - /.22 | - /.19 | - /1.1 | - /.19 | - /.16 | - /1.1 | - /.17 |
| No High. | - | .19 | - | - | - | - | - | - | - | |
| Oracle | .19 | .12 | .33/.39 | .07/.15 | 4.7/2.6 | .54/.39 | .15/.15 | 3.6/2.6 | .41/.39 | |
| Unsup. | .15 | .13 | .25/.33 | .11/.17 | 2.2/1.9 | .30/.26 | .13/.12 | 2.3/2.1 | .27/.29 | |
| Sup. | .12 | .10 | .30/.36 | .08/.17 | 3.7/2.1 | .36/.23 | .10/.10 | 3.6/2.3 | .33/.28 | |
| Domain | Speed | |||||||||
| EnglishItalian | ||||||||||
| Biomed. | Fast | .09 | .04/.01 | .12/.02 | .03/.01 | 4.0/2.0 | .30/.27 | .08/.11 | 3.7/2.4 | .17/.04 |
| Avg. | .10/.05 | .27/.12 | .09/.04 | 3.0/3.0 | .22/.30 | .07/.11 | 3.1/2.7 | .24/.17 | ||
| Slow | .09/.02 | .21/.04 | .08/.01 | 2.6/4.0 | .19/.26 | .07/.11 | 2.7/2.3 | .20/.07 | ||
| Social | Fast | .14 | .11/.07 | .30/.20 | .07/.04 | 4.2/5.0 | .40/.52 | .11/.16 | 3.6/3.2 | .34/.29 |
| Avg. | .23/.14 | .48/.32 | .18/.10 | 2.6/3.2 | .30/.42 | .09/.15 | 3.3/2.8 | .37/.36 | ||
| Slow | .17/.05 | .39/.14 | .14/.03 | 2.7/4.6 | .31/.54 | .11/.17 | 2.8/3.1 | .35/.22 | ||
| EnglishDutch | ||||||||||
| Biomed. | Fast | .14 | .03/.02 | .11/.05 | .02/.01 | 5.5/5.0 | .48/.61 | .13/.18 | 3.6/3.3 | .18/.09 |
| Avg. | .11/.19 | .20/.30 | .10/.17 | 2.0/1.7 | .25/.29 | .13/.16 | 1.9/1.8 | .22/.29 | ||
| Slow | .12/.10 | .26/.23 | .10/.07 | 2.6/3.2 | .29/.42 | .12/.16 | 2.4/2.6 | .27/.30 | ||
| Social | Fast | .12 | .06/.07 | .19/.21 | .04/.04 | 4.7/5.2 | .37/.47 | .10/.13 | 3.7/3.6 | .25/.29 |
| Avg. | .17/.32 | .32/.48 | .15/.29 | 2.1/1.6 | .22/.23 | .10/.12 | 2.2/1.9 | .26/.31 | ||
| Slow | .18/.18 | .38/.40 | .15/.14 | 2.5/2.8 | .25/.34 | .09/.11 | 2.7/3.0 | .30/.37 | ||
| Language | MQM Category | MT | No Highlight | Oracle | Unsupervised | Supervised | |||||
| Maj. | Min. | Maj. | Min. | Maj. | Min. | Maj. | Min. | Maj. | Min. | ||
| Italian | Accuracy - Addition | 0 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 1 |
| Accuracy - Mistranslation | 21 | 22 | 10 | 12 | 4 | 8 | 24 | 17 | 17 | 17 | |
| Accuracy - Inconsistency | 2 | 4 | 1 | 3 | 2 | 2 | 1 | 3 | 0 | 2 | |
| Accuracy - Omission | 2 | 0 | 0 | 0 | 0 | 1 | 4 | 1 | 1 | 2 | |
| Accuracy - Untranslated | 1 | 4 | 1 | 2 | 0 | 1 | 1 | 1 | 3 | 2 | |
| Style - Inconsistent Style | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |
| Style - Readability | 17 | 25 | 5 | 30 | 0 | 12 | 4 | 34 | 1 | 29 | |
| Style - Wrong Register | 0 | 8 | 0 | 3 | 0 | 3 | 1 | 1 | 3 | 2 | |
| Linguistic - Grammar | 6 | 15 | 2 | 16 | 0 | 5 | 3 | 12 | 2 | 12 | |
| Linguistic - Punctuation | 1 | 13 | 0 | 9 | 0 | 3 | 1 | 6 | 0 | 3 | |
| Linguistic - Spelling | 5 | 3 | 0 | 4 | 0 | 3 | 3 | 2 | 0 | 1 | |
| Total | 55 | 95 | 19 | 79 | 6 | 38 | 42 | 77 | 28 | 71 | |
| Dutch | Accuracy - Addition | 0 | 1 | 0 | 2 | 0 | 3 | 0 | 2 | 0 | 1 |
| Accuracy - Mistranslation | 25 | 34 | 18 | 25 | 23 | 27 | 12 | 31 | 16 | 29 | |
| Accuracy - Inconsistency | 0 | 0 | 0 | 2 | 0 | 2 | 0 | 2 | 0 | 5 | |
| Accuracy - Omission | 3 | 1 | 1 | 1 | 2 | 1 | 1 | 1 | 4 | 2 | |
| Accuracy - Untranslated | 4 | 4 | 1 | 1 | 1 | 4 | 1 | 3 | 0 | 2 | |
| Style - Inconsistent Style | 2 | 0 | 0 | 5 | 1 | 7 | 0 | 2 | 0 | 9 | |
| Style - Readability | 1 | 27 | 1 | 20 | 0 | 13 | 2 | 15 | 6 | 41 | |
| Style - Wrong Register | 0 | 2 | 0 | 3 | 0 | 3 | 0 | 1 | 1 | 0 | |
| Linguistic - Grammar | 3 | 19 | 2 | 14 | 3 | 23 | 2 | 6 | 3 | 12 | |
| Linguistic - Punctuation | 0 | 6 | 0 | 3 | 0 | 4 | 0 | 2 | 0 | 3 | |
| Linguistic - Spelling | 1 | 1 | 1 | 1 | 2 | 1 | 0 | 1 | 0 | 0 | |
| Total | 39 | 95 | 24 | 77 | 32 | 88 | 18 | 66 | 30 | 104 | |
| Method | DivEMT | QE4PE | ||||||
| EnIt | EnNl | EnIt | EnNl | |||||
| AP | AU | AP | AU | AP | AU | AP | AU | |
| Logprobs (Fomicheva et al., 2020) | 0.18 | 0.18 | 0.19 | 0.19 | 0.10 | 0.09 | 0.09 | 0.09 |
| Prediction Depth (Baldock et al., 2021) | 0.26 | 0.50 | 0.26 | 0.51 | 0.17 | 0.23 | 0.19 | 0.27 |
| Logprobs (Fomicheva et al., 2020, Unsup.) | 0.41 | 0.41 | 0.42 | 0.42 | 0.23 | 0.23 | 0.31 | 0.31 |
| XCOMET-XXL (Guerreiro et al., 2024b, Sup.) | 0.16 | 0.23 | 0.19 | 0.28 | ||||
| avg. Oracle single translator | - | - | - | - | 0.53 | 0.73 | 0.55 | 0.75 |
Pre-task
Main task
Post-task
Pre-task
Main task
Post-task