### v2

# QE4PE: Word-level Quality Estimation for Human Post-Editing

Abstract
Word-level quality estimation (QE) methods aim to detect erroneous spans in machine translations, which can direct and facilitate human post-editing. While the accuracy of word-level QE systems has been assessed extensively, their usability and downstream influence on the speed, quality and editing choices of human post-editing remain understudied. In this study, we investigate the impact of word-level QE on machine translation (MT) post-editing in a realistic setting involving 42 professional post-editors across two translation directions. We compare four error-span highlight modalities, including supervised and uncertainty-based word-level QE methods, for identifying potential errors in the outputs of a state-of-the-art neural MT model. Post-editing effort and productivity are estimated from behavioral logs, while quality improvements are assessed by word- and segment-level human annotation. We find that domain, language and editors’ speed are critical factors in determining highlights’ effectiveness, with modest differences between human-made and automated QE highlights underlining a gap between accuracy and usability in professional workflows.
1 Introduction
Recent years have seen a steady increase in the quality of machine translation (MT) systems and their widespread adoption in professional translation workflows (Kocmi et al., 2024a). Still, human post-editing of MT outputs remains a fundamental step to ensure high-quality translations, particularly for challenging textual domains requiring native fluency and specialized terminology (Liu et al., 2024). Quality estimation (QE) techniques were introduced to reduce post-editing effort by automatically identifying problematic MT outputs without the need for human-written reference translations and were quickly integrated into industry platforms (Tamchyna, 2021). Segment-level QE models correlate well with human perception of quality (Freitag et al., 2024) and exceed the performance of reference-based metrics in specific settings (Rei et al., 2021; Amrhein et al., 2022, 2023). On the other hand, word-level QE methods for identifying error spans requiring revision have received less attention in the past due to their modest agreement with human annotations, despite their promise for more granular and interpretable quality assessment in line with modern MT practices (Zerva et al., 2024). In particular, while the accuracy of these approaches is regularly assessed in evaluation campaigns, research has rarely focused on assessing the impact of such techniques in realistic post-editing workflows, with notable exceptions suggesting limited benefits (Shenoy et al., 2021; Eo et al., 2022). This hinders current QE evaluation practices: by foregoing experimental evaluation with human editors, it is implicitly assumed that word-level QE will become helpful once sufficient accuracy is reached, without accounting for the additional challenges towards a successful integration of these methods in post-editing workflows.
In this study, which we dub QE4PE (Quality Estimation for Post Editing), we address this gap by conducting a large-scale study with 42 professional translators for the EnglishItalian and EnglishDutch directions to measure the impact of word-level QE on editing quality, productivity and usability. We aim for a realistic and reproducible setup, employing the high-quality open-source NLLB 3.3B MT model (NLLB Team et al., 2024) to translate challenging documents from biomedical and social media domains. We then conduct a controlled evaluation of post-editing with error spans in four highlight modalities, i.e. using highlights derived from four word-level QE methods: a supervised state-of-the-art QE model trained on human error annotations (XCOMET, Guerreiro et al., 2024), an unsupervised method leveraging the uncertainty of the MT model during generation, oracle error spans obtained from the consensus of previous human post-editors, and a no highlight baseline. The human post-editing is performed using GroTE, a simple online interface we built to support the real-time logging of granular editing data, enabling a quantitative assessment of editing effort and productivity across highlight modalities. We also survey professionals using an online questionnaire to collect qualitative feedback about the usability and quality of the MT model, as well as the interface and error span highlights. Finally, a subset of the original MT outputs and their post-edited variants is annotated following the MQM and ESA protocols (Lommel et al., 2013; Kocmi et al., 2024b) to verify quality improvements after post-editing. See Figure˜1 for an overview of the study. Our work represents a step towards an evaluation of translation technologies that is centered on users’ experience (Guerberof-Arenas and Moorkens, 2023; Savoldi et al., 2025).
We release all data, code and the GroTE editing interface to foster future studies on the usability of error span highlighting techniques for other word-level QE methods and translation directions.111Data: hf.co/gsarti/qe4pe. Code: gsarti/qe4pe.
2 Related Work
MT Post-editing
Human post-editing of MT outputs is increasingly common in professional translator workflows, as it was shown to increase the productivity of translators while preserving translation quality across multiple domains (Liu et al., 2024). However, many factors were found to influence the variability of post-editing productivity across setups, including MT quality (Zouhar et al., 2021b), interface familiarity (Läubli et al., 2022), individual variability and source-target languages typological similarity (Sarti et al., 2022). Studies evaluating the post-editing process generally focus on productivity, i.e. number of processed words/characters per minute, and the temporal, technical and cognitive dimensions of post-editing effort, operationalized through behavioral metrics such as editing time, keystrokes and pauses (Krings, 2001; Sarti et al., 2022). We adopt these metrics for the QE4PE study and relate them to different highlight modalities.
Quality estimation for MT
The field of quality estimation was initially concerned with MT model uncertainty (Blatz et al., 2004; Specia et al., 2009), but in time began focusing on predicting translation quality even without using references (Turchi et al., 2013, 2014; Kepler et al., 2019; Thompson and Post, 2020 inter alia). Advances in segment- and word-level QE research are regularly assessed in annual WMT campaigns (Fomicheva et al., 2021; Zerva et al., 2022, 2024; Blain et al., 2023), where the best-performing QE systems are usually Transformer-based language models trained on human quality judgments, such as the popular COMET model suite (Rei et al., 2020, 2021, 2022). The widespread adoption of the fine-grained Multidimensional Quality Metrics scale (MQM, Lommel et al., 2013) prompted a paradigm shift in MT evaluation, leading to new QE metrics predicting quality at various granularity levels (Kocmi and Federmann, 2023; Fernandes et al., 2023; Guerreiro et al., 2024). Aside from supervised models, unsupervised methods exploiting model uncertainty and its internal mechanisms were proposed as efficient alternatives to identify potential error spans in MT outputs (Fomicheva et al., 2020; Dale et al., 2023; Xu et al., 2023; Himmi et al., 2024, surveyed by Leiter et al. 2024). In this work, we compare the downstream effectiveness of state-of-the-art supervised and unsupervised word-level QE metrics for post-editing settings.
QE for Human Post-Editing Workflows
Automatic QE methods are widely used in the translation industry for triaging automatic translations (Tamchyna, 2021). While QE usage has been found helpful to increase the confidence and speed of human assessment (Mehandru et al., 2023; Zouhar et al., 2025), an incautious usage of these techniques can lead to a misplaced over-reliance on model predictions (Zouhar et al., 2021a). Interfaces supporting word-level error highlights were developed for studying MT post-editing (Coppers et al., 2018; Herbig et al., 2020) and code reviewing (Sun et al., 2022; Vasconcelos et al., 2025), with results suggesting that striking the right balance of user-provided information is fundamental to improve the editing experience and prevent cognitive overload. Most similar to our study, Shenoy et al. (2021) investigated the effect of synthetic word-level QE highlights for EnglishGerman post-editing on Wikipedia data, concluding that word-level QE accuracy was at the time still insufficient to produce tangible productivity benefits in human editing workflows. In this work, we expand the scope of such evaluation by including two translation directions, two challenging real-world text domains and state-of-the-art MT and QE systems and methods.
3 Experimental Setup
3.1 Structure of the Study
Our study is organized in five stages:
1) Oracle Post-editing
As a preliminary step, segments later used in the main assessment are post-edited by three professionals per direction using their preferred interface without logging. This allows us to obtain post-edits and produce oracle word-level spans based on the editing consensus of multiple human professionals. Translators involved in this stage are not involved further in the study.
2) Pretask (Pre)
The pretask allows the core translators (12 per language direction, see Section˜3.4) to familiarize themselves with the GroTE interface and text highlights. Before starting, all translators complete a questionnaire to provide demographic and professional information about their profile (Table˜10). In the pretask, all translators work in an identical setup, post-editing a small set of documents similar to those of the main task with supervised highlights. We assign core translators into three groups based on their speed from editing logs (4 translators per group for faster, average and slower groups in each direction). Individuals from each group are then assigned randomly to each highlight modality to ensure an equal representation of editing speeds, resulting in 1 faster, 1 average and 1 slower translator for each highlight modality. This procedure is repeated independently for both translation directions.
3) Main Task (Main)
This task, conducted in the two weeks following the pretask, covers the majority of the collected data and is the main object of study for the analyses of Section˜4. In the main task, 24 core translators work on the same texts using the GroTE interface, with three translators per modality in each translation direction, as shown in Figure˜1. After the main task, translators complete a questionnaire on the quality and usability of the MT outputs, the interface and, where applicable, word highlights.222We do not disclose the highlight modality to translators to avoid biasing their judgment in the evaluation.
4) Post-Task (Post)
After Main, the 12 core translators per direction are asked to post-edit an additional small set of related documents with GroTE, but this time working all with the no highlight modality. This step lets us obtain baseline editing patterns for each translator to estimate individual speed and editing differences across highlight modalities without the confounder of interface proficiency accounted for in the Pre stage.
5) Quality Assessment (QA)
Finally, a subset consisting of 148 main task segments is randomly selected for manual annotation by six new translators per direction (see Section˜3.4). For each segment, the original MT output and all its post-edited versions are annotated with MQM error spans, including minor/major error severity and a subset of MQM error categories including e.g. mistranslations, omissions and stylistic errors (Lommel et al., 2013).333See Figure 5 for an overview of setup and guidelines. Moreover, the annotator proposes corrections for each error span, ultimately providing a 0-100 quality score matching the common DA scoring adopted in multiple WMT campaigns. We adopt this scoring system, which closely adheres to the ESA evaluation protocol (Kocmi et al., 2024b), following recent results showing its effectiveness and efficiency for ranking MT system.
In summary, for each translation direction, we collect 3 full sets of oracle post-edits, 12 full sets of edits with behavioral logs for pre, main and post task data, and 13 subsets of main task data (12 post-edits, plus the original MT output) annotated with MQM error spans, corrections and segment-level ESA ratings. Moreover, we also collect 12 pre- and post-task questionnaire responses from core set translators to obtain a qualitative view of the editing process.
3.2 Highlight Modalities
We conduct our study on four highlight modalities across two severity levels (minor and major errors). Using multiple severity levels follows the current MT evaluation practices (Freitag et al., 2021, 2024), and previous results showing that users tend to prefer more granular and informative word-level highlights (Shenoy et al., 2021; Vasconcelos et al., 2025). The highlight modalities we employ are:
No Highlight
The text is presented as-is, without any highlighted spans. This setting serves as a baseline to estimate the default post-editing quality and productivity using our interface.
Oracle
Following the Oracle Post-editing phase, we produce oracle error spans from the editing consensus of human post-editors. We label text spans that were edited by two out of three translators as minor, and those edited by all three translators as major, following the intuition that more critical errors are more likely to be identified by several annotators, while minor changes will show more variance across subjects. This modality serves as a best-case scenario, providing an upper bound for future improvements in word-level QE quality.
Supervised
In this setting, word-level error spans are obtained using XCOMET-XXL (Guerreiro et al., 2024), which is a multilingual Transformer encoder (Goyal et al., 2021) further trained for joint word- and sentence-level QE prediction. We select XCOMET-XXL in light of its broad adoption, open accessibility and state-of-the-art performance in QE across several translation directions (Zerva et al., 2024). For the severity levels, we use the labels predicted by the model, mapping critical labels to the major level.
| Method | DivEMT | QE4PE | ||
| en-it | en-nl | en-it | en-nl | |
| Logprobs | 0.18 | 0.19 | 0.10 | 0.09 |
| MCD Var. | 0.41 | 0.42 | 0.23 | 0.31 |
| XCOMET (Sup.) | 0.34 | 0.35 | 0.16 | 0.19 |
| Avg. Trans. | - | - | 0.53 | 0.55 |
Unsupervised
In this modality, we exploit the access to the MT model producing the original translations to obtain uncertainty-based highlights. As a preliminary evaluation to select a capable unsupervised word-level QE method, we evaluate two unsupervised QE methods employing token log-probabilities assigned by MT model to predict human post-edits: raw negative log-probabilities (Logprobs), corresponding to the surprisal assigned by the MT model to every generated token, and their variance for 10 steps of Monte Carlo Dropout (MCD Var., Gal and Ghahramani, 2016). We employ surprisal-based metrics following previous work showing their effectiveness in predicting translation errors (Fomicheva and Specia, 2019) and human editing time (Lim et al., 2024). We collect scores for the EnglishItalian and EnglishDutch directions of QE4PE Oracle post-edits and DivEMT (Sarti et al., 2022) to identify the best-performing method, using metric scores extracted from the original models used for translation to predict human post-edits. We use average precision (AP) as a threshold-agnostic performance metric for the tested continuous methods. Oracle highlights obtained from the consensus of three annotator in the first stage of the study are used as reference for QE4PE, while a single set of post-edits is available for DivEMT. The XCOMET-XXL model used for Supervised highlights, and the average agreement of individual Oracle editors with the consensus label are also included for comparison. Table˜1444Full results in Table 16. Highlights are extended from tokens to words to match the granularity of other modalities. show a strong performance for the MCD Var. method, even surpassing the accuracy of the supervised XCOMET model across both datasets. Hence, we select MCD Var. for the Unsupervised highlight modality, setting value thresholds for minor/major errors to match the respective highlighted word proportions in the Supervised modality to ensure a fair comparison.
3.3 Data and MT model
MT Model
On the one hand, the MT model must achieve high translation quality in the selected languages to ensure our experimental setup applies to state-of-the-art proprietary systems. Still, the MT model should be open-source and have a manageable size to ensure reproducible findings and enable the computation of uncertainty for the unsupervised setting. All considered, we use NLLB 3.3B (NLLB Team et al., 2024), a widely-used MT model achieving industry-level performances across 200 languages (Moslem et al., 2023).
| Task | Domain | # Docs | # Seg. | # Words |
| Pre | Social | 4 | 23 | 539 |
| Biomed. | 2 | 15 | 348 | |
| Main | Social | 30 | 160 | 3375 |
| Biomed. | 21 | 165 | 3384 | |
| Post | Social | 6 | 34 | 841 |
| Biomed. | 2 | 16 | 257 | |
| Total | 64 | 413 | 8744 | |
Data selection
We begin by selecting two translation directions, EnglishItalian and EnglishDutch, according to the availability of professional translators from our industrial partners. We intentionally focus on out-of-English translations as they are generally more challenging for modern MT models (Kocmi et al., 2023). We aim to identify documents that are manageable for professional translators without domain-specific expertise but still prove challenging for our MT model to ensure a sufficient amount of error spans across modalities. Since original references for our selected translation direction were not available, we do not have a direct mean to compare MT quality in the two languages. However, according to our human MQM assessment in Section˜4.3 (Table˜5), NLLB produces a comparable amount of errors across Dutch and Italian MT, suggesting similar quality.
We begin by translating 3,672 multi-segment English documents from the WMT23 General and Biomedical MT shared tasks (Kocmi et al., 2023; Neves et al., 2023) and MT test suites to Dutch and Italian. Our choice for these specialized domains, as opposed to e.g. generic news articles, is driven by the real-world needs of the translation industry for domain-specific post-editing support (Eschbach-Dymanus et al., 2024; Li et al., 2025). Moreover, focusing on domains that are considerably more challenging for MT systems than news, as shown by recent WMT campaigns (Neves et al., 2024), ensures a sufficient amount of MT errors to support a sound comparison of word-level QE methods. Then, XCOMET-XXL is used to produce a first set of segment-level QE scores and word-level error spans for all segments. To make the study tractable, we further narrow down the selection of documents according to several heuristics to ensure a realistic editing experience and a balanced occurrence of error spans (details in Appendix˜A). This procedure yields 351 documents, from which we manually select a subset of 64 documents (413 segments, 8,744 source words per post-editor) across two domains:
| Source en | So why is it that people jump through extra hoops to install Google Maps? |
| No High. | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| Oracle | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| Sup. | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| Unsup. | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| PENo High. | Quindi perché le persone devono fare un passaggio in più per installare Google Maps? |
| PEOracle | Allora, perché le persone fanno un passaggio in più per installare Google Maps? |
| PESup. | Quindi perché le persone fanno passaggi in più per installare Google Maps? |
| PEUnsup. | Quindi perché le persone fanno i salti mortali per installare Google Maps? |
Biomedical abstracts extracted from PubMed from the WMT23 Biomedical Translation Task (Neves et al., 2023), including domain-specific terminology.
Table˜2 present statistics for the Pre, Main and Post editing stages, and Table˜3 shows an example of highlights and edits. While the presence of multiple domains in the same task can render our post-editing setup less realistic, we deem it essential to test the cross-domain validity of our findings.
Critical Errors
Before producing highlights, we manually introduce 13 critical errors in main task segments to assess post-editing thoroughness. Errors are produced, for example, by negating statements, inverting the polarity of adjectives, inverting numbers, and corrupting acronyms. We replicate the errors in both translation directions to enable direct comparison. Most of these errors were correctly identified across all three highlight modalities (examples in Table˜7).
3.4 Participants
For both directions, professional translation companies Translated Srl555https://translated.com and Global Textware666https://globaltextware.nl/ recruited three translators for the Oracle post-editing stage, the core set of 12 translators working on Pre, Main and Post tasks, and six more translators for the QA stage, for a total of 21 translators per direction. All translators were freelancers with native proficiency in their target language and self-assessed proficiency of at least C1 in English. Almost all translators had more than two years of professional translation experience and regularly post-edited MT outputs (details in Table˜10).
3.5 Editing Interface
We develop a custom interface, which we name Groningen Translation Environment (GroTE, Figure˜2), to support editing over texts with word-level highlights. While the MMPE tool used by Shenoy et al. (2021) provide extensive multimodal functionalities (Herbig et al., 2020), we aim for a bare-bones setup to avoid confounders in the evaluation. GroTE is a web interface based on Gradio (Abid et al., 2019) and hosted on the Hugging Face Spaces to enable multi-user data collection online. Upon loading a document, source texts and MT outputs for all segments are presented in two columns following standard industry practices. For modalities with highlights, the interface provides an informative message and supports the removal of all highlights on a segment via a button, with highlights on words disappearing automatically upon editing, as in Shenoy et al. (2021). The interface supports real-time logging of user actions, allowing for the analysis of the editing process. In particular, we log the start/end times for each document, the accessing and exiting of segment textboxes, highlight removals, and keystrokes.
GroTE intentionally lacks standard features such as translation memories, glossaries, and spellchecking to ensure equal familiarity among translators, ultimately controlling for editor proficiency with these tools, as done in previous studies (Shenoy et al., 2021; Sarti et al., 2022). While most translators noted the lack of features in our usability assessment, the majority also found the interface easy to set up, access, and use (Table˜10).
4 Analysis
4.1 Productivity
We obtain segment- and document-level edit times and compute editing productivity as the number of processed source characters over the sum of all document-level edit times, measured in characters per minute. To account for potential breaks taken by post-editors during editing, we filter out pauses between logged actions longer than 5 minutes. We note that this procedure does not significantly impact the overall ranking of translators, while ensuring a more robust evaluation of editing time.
Do Highlights Make Post-editors Faster?
Figure˜3 shows translators’ productivity across stages, with every dot corresponding to the productivity of a single individual. We observe that no highlight modality leads to systematically faster editing across all speed groups and that the ordering of Pre-task speed groups is maintained in the following stages despite the different highlight modalities. These results suggest that individual variability in editing speed is more critical than highlight modality in predicting editing speed. However, faster EnglishDutch translators achieve outstanding productivity, i.e. standard deviations above the overall mean ( char/min, ➔ in Figure˜3) almost exclusively in No Highlight, and, Oracle modalities, suggesting that lower-quality highlights hinder editing speed.
We validate these observations by fitting a negative binomial mixed-effect model on segment-level editing times (model details in Table˜8). Excluding random factors such as translator and segment identity from the model produces a significant drop in explained variance, confirming the inherent variability of editing times (). Model coefficients show that MT output length and the proportion of highlighted characters are the main factors driving an increase in editing times, possibly reflecting an increase in cognitive effort to process additional information. We find highlights to have a significant impact on increasing the editing speed of EnglishItalian translators , but a minimal impact for EnglishDutch. Comparing the productivity of the same translator editing with and without highlights (Main vs Post), two-thirds of the translators editing with highlights were up to two times slower on biomedical texts. However, the same proportion of translators was up to three times faster on social media texts across both directions.
In summary, we find that highlight modalities are not predictive of edit times on their own, but translation direction and domain play an important role in determining the effect of highlights on editing productivity. We attribute these results to two main factors, which will remain central in the analysis of the following sections: (1) the different propensity of translators to act upon highlighted issues in the two tested directions, and (2) the different nature of errors highlighted across domains.
4.2 Highlights and Edits
We then examine how highlights are distributed across modalities and how they influence the editing choices of human post-editors.
| Base Freq. | Measured | Projected | ||||||||
| EnglishItalian | ||||||||||
| No High. | - | 0.05 | - | - | - | - | - | - | - | - |
| Random | 0.16 | - | - | - | - | - | 0.06 | 1.20 | 0.18 | 1.20 |
| Oracle | 0.15 | 0.12 | 0.37 | 4.62 | 0.45 | 4.1 | 0.18↓0.19 | 6.00↑1.38 | 0.55↑0.10 | 4.23↑0.14 |
| Unsup. | 0.16 | 0.13 | 0.25 | 2.27 | 0.21 | 2.2 | 0.11↓0.14 | 2.75↑0.48 | 0.37↑0.16 | 2.47↑0.26 |
| Sup. | 0.12 | 0.16 | 0.28 | 2.00 | 0.22 | 2.0 | 0.14↓0.14 | 3.50↑1.50 | 0.35↑0.13 | 3.18↑1.18 |
| EnglishDutch | ||||||||||
| No High. | - | 0.14 | - | - | - | - | - | - | - | - |
| Random | 0.17 | - | - | - | - | - | 0.16 | 1.14 | 0.19 | 1.19 |
| Oracle | 0.20 | 0.10 | 0.26 | 4.33 | 0.53 | 3.12 | 0.28↑0.02 | 2.55↓1.78 | 0.40↓0.13 | 2.35↓0.77 |
| Unsup. | 0.20 | 0.11 | 0.20 | 2.50 | 0.36 | 2.00 | 0.22↑0.02 | 1.83↓0.67 | 0.31↓0.05 | 1.72↓0.28 |
| Sup. | 0.12 | 0.09 | 0.24 | 3.43 | 0.33 | 3.30 | 0.28↑0.04 | 2.33↓1.10 | 0.24↓0.09 | 2.40↓0.90 |
Agreement Across Modalities
First, we quantify how different modalities agree in terms of highlight distribution and editing. We find that highlight overlaps across modalities range between 15% and 39% when comparing highlight modalities in a pairwise fashion, with the highest overlap for EnglishItalian social media and EnglishDutch biomedical texts.777Scores are normalized to account for highlight frequencies across modalities. Agreement is shown in Table 11. Despite the relatively low highlight agreement, we find an average agreement of 73% for post-edited characters across modalities. This suggests that edits are generally uniform regardless of highlight modalities and are not necessarily restricted to highlighted spans.888Editing agreement is shown in Figure 7.
Do Highlights Accurately Identify Potential Issues?
Table˜4 (Base Freq.) shows raw highlight and edit frequencies across modalities. We observe different trends across the two language pairs: for EnglishItalian, post-editors working with highlights edit more than twice as much as translators with No Highlight, regardless of the highlight modality. On the contrary, for EnglishDutch they edit 33% less in the same setting. These results suggest a different attitude towards acting upon highlighted potential issues across the two translation directions, with EnglishItalian translators appearing to be conditioned to edit more when highlights are present. We introduce four metrics to quantify highlights-edits overlap:
and , reflecting highlights’ precision and recall in predicting edits, respectively.
shows how much more likely an edit is to fall within rather than outside highlighted characters.
shows how much more likely it is for a highlight to mark edited rather than unmodified spans.
Intuitively, character-level recall should be more indicative of highlight quality compared to precision , provided that word-level highlights can be useful even when not minimal.999For example, if the fully-highlighted word traduttore is changed to its feminine version traduttrice, (edit correctly and fully predicted) but since word stem characters are left unchanged. Table˜4 (Measured) shows metric values across the three highlight modalities (breakdowns by domain and speed shown in Tables˜13 and 14). As expected, Oracle highlights obtain the best performance in terms of precision and recall, with , in particular, being significantly higher than the other two modalities across both directions.
Surprisingly, we find no significant precision and recall differences between Supervised and Unsupervised highlights, despite the word-level QE training of XCOMET used in the former modality. Moreover, they support the potential of unsupervised, model internals-based techniques to complement or substitute more expensive supervised approaches. Still, likelihood ratios for all modalities and directions indicate that highlights are 2-4 times more likely to precisely and comprehensively encompass edits than non-highlighted texts. This suggests that even imperfect highlights that do not reach Oracle-level quality might effectively direct editing efforts toward potential issues. We validate these observations by fitting a zero-inflated negative binomial mixed-effects model to predict segment-level edit rates. Results confirm a significantly higher edit rate for EnglishItalian highlighted modalities and the social media domain with (features and significances shown in Appendix Table˜9). We find a significant zero inflation associated with translator identity, suggesting the choice of leaving MT outputs unedited is highly subjective.
Do Highlights Influence Editing Choices?
Since in Section˜4.1 we found the proportion of highlighted characters to impact the editing rate of translators, we question whether the relatively high and values might be artificially inflated by the eagerness of translators to intervene on highlighted spans. In other words, do highlights identify actual issues, or do they condition translators to edit when they otherwise would not? To answer this, we propose to project highlights from a selected modality—in which highlights were shown during editing—onto the edits performed by the No Highlight translators on the same segments. The resulting difference between measured and projected metrics can then be taken as an estimate for the impact of highlight presentation on their resulting accuracy.
To further ensure the soundness of our analysis, we use a set of projected Random highlights as a lower bound for highlight performance. To make the comparison fair, Random highlights are created by randomly highlighting words in MT outputs matching the average word-level highlight frequency across all highlighted modalities given the current domain and translation direction. Table˜4 (Projected) shows results for the three highlighted modalities. First, all projected metrics remain consistently above the Random baseline, suggesting a higher-than-chance ability to identify errors even for worst-performing highlight modalities. Projected precision scores depend on edit frequency, and hence see a major decrease for EnglishItalian, where the No Highlight edit rate is much lower. However, the increase in across all EnglishItalian modalities confirms that, despite the lower edit proportion, highlighted texts remain notably more likely to be edited than non-highlighted ones. Conversely, the lower , and for EnglishDutch show that edits become much less skewed towards highlighted spans in this direction when accounting for presentation bias.
Overall, while the presence of highlights makes EnglishItalian translators more likely to intervene in MT outputs, their location in the MT output often pinpoints issues that would be edited regardless of highlighting. EnglishDutch translators, on the contrary, intervene at roughly the same rate regardless of highlights presence, but their edits are focused mainly on highlighted spans when they are present. This difference is consistent across all subjects in the two directions despite the identical setup and comparable MT and QE quality across languages. This suggests that cultural factors might play a non-trivial role in determining the usability and influence of QE methods regardless of span accuracy, a phenomenon previously observed in human-AI interaction studies (Ge et al., 2024).
4.3 Quality Assessment
We continue our assessment by inspecting the quality of MT and post-edited outputs along three dimensions. First, we use XCOMET segment-level QE ratings as an automatic approximation of quality and compare them to human-annotated quality scores collected in the last phase of our study. For efficiency, these are obtained for the 0-100 Direct Assessment scale commonly used in QE evaluation (Specia et al., 2020), but following an initial step of MQM error annotation to condition scoring on found errors, as prescribed by the ESA protocol (Kocmi et al., 2024b). Then, MQM error span annotations are used to analyze the distribution of error categories. Finally, we manually assess critical errors, which were inserted to quantify highlight modalities effect on unambiguous issues.
Do Highlights Influence Post-Editing Quality?
In this stage, we focus particularly on edited quality improvements, i.e. how post-editing the same MT outputs under different highlight conditions influences the resulting quality of translations. We operationalize this assessment using human ratings and automatic metrics to score MT and post-edited translations, using their difference as the effective quality gain after the post-editing stage. Scores for this metric are generally positive, i.e. human post-editing improves quality, and bounded by the maximal achievable quality gain given the initial MT quality. Figure˜4 shows median improvement values across quality bins defined from the distribution of initial MT quality scores (shown in histograms), in which all post-edited versions of each MT output appear as separate observations. Positive median scores confirm that post-edits generally lead to quality improvements across all tested settings. However, we observe different trends across the two metrics: across both domains, XCOMET greatly underestimates the human-assessed ESA quality improvement, especially for biomedical texts where it shows negligible improvement regardless of the initial MT quality. These results echo recent findings cautioning users against the poor performance of trained MT metrics for unseen domains and high-quality translations (Agrawal et al., 2024; Zouhar et al., 2024). Focusing on the more reliable ESA scores, we observe large quality improvements from post-editing, as shown by near-maximal quality gains across most bins and highlight modalities. While No Highlight seems to underperform other modalities in the social media domain, the lack of more notable differences in gains across highlight modalities suggests that highlights’ quality impact might not be evident in terms of segment-level quality, motivating our next steps in the quality analysis.
We also find no clear relationship between translator speed and edited quality improvements, suggesting that higher productivity does not come at a cost for faster translators (Figure˜10). This finding confirms that neglecting errors is not the cause of different editing patterns from previous sections.
| MT | No High. | Oracle | Unsup. | Sup. | ||
| Italian | Acc. | 26 / 31 | 12 / 17 | 6 / 12 | 30 / 22 | 22 / 24 |
| Style | 17 / 33 | 5 / 33 | 0 / 15 | 5 / 35 | 4 / 31 | |
| Ling. | 12 / 31 | 2 / 29 | 0 / 11 | 7 / 20 | 2 / 16 | |
| Tot. | 55 / 95 | 19 / 79 | 6 / 38 | 42 / 77 | 28 / 71 | |
| Dutch | Acc. | 32 / 40 | 20 / 31 | 26 / 37 | 14 / 39 | 20 / 39 |
| Style | 3 / 29 | 1 / 28 | 1 / 23 | 2 / 18 | 7 / 50 | |
| Ling. | 4 / 26 | 3 / 18 | 5 / 28 | 2 / 9 | 3 / 15 | |
| Tot. | 39 / 95 | 24 / 77 | 32 / 88 | 18 / 66 | 30 / 104 |
Which Error Types Do Highlights Identify?
Table˜5 shows a breakdown of MQM annotations for MT and all highlight modalities using the Accuracy, Style and Linguistic macro-categories of MQM errors.101010Full micro-category breakdown in Table 12, per-domain breakdown in Figure 9. Category descriptions in Figure 5. At this granularity, differences across modalities become visible, with overall error counts showing a clear relation to from Table˜4 (Oracle being remarkably better for EnglishItalian, with milder and more uniform trends in EnglishDutch). At least for EnglishItalian, these results confirm that an observable quality improvement from editing with highlights is present in the best-case Oracle scenario. By contrast, for EnglishDutch the Unsupervised method is found to outperform even the Oracle setting in reducing the amount of errors, while it fares relatively poorly for EnglishItalian. We also note a different distribution of Accuracy and Style errors, with the former being more common in biomedical texts while the latter appearing more often for translated social media posts (Figure˜9). We posit that differences in error types across domains might explain the opposite productivity trends observed in Section˜4.1: while highlighted accuracy errors might lead to time-consuming terminology verification in biomedical texts, style errors might be corrected more quickly and naturally in the social media domain.
Do Highlights Detect Critical Errors?
We examine whether the critical errors we inserted were detected by different modalities, finding that while most modalities fare decently with more than 62% of critical errors highlighted, Unsupervised is the only setting for which all errors are correctly highlighted across both directions. Then, critical errors are manually verified in all outputs, finding that 16-20% more critical errors are edited in highlighted modalities compared to No Highlight (full results in Table˜12). Hence, highlights might lead to narrow but tangible quality improvements that can go undetected in coarse quality assessments, and finer-grained evaluations might be needed to quantify future improvements in word-level QE.
4.4 Usability
In post-task questionnaire answers (Table˜6), most translators stated that MT outputs had average-to-high quality and that provided texts were challenging to translate. Highlights were generally found decently accurate, but they were generally not found useful to improve either productivity or quality (including Oracle ones). Interestingly, despite the convincing gains for critical errors measured in the last section, most translators stated that highlights did not influence their editing and did not help them identify errors that would have otherwise been missed. Concretely, this suggests that the potential quality improvements might not be easily perceived by translators and might have secondary importance compared to the extra cognitive load elicited by highlighted spans. When asked to comment about highlights, several translators called them “more of an eye distraction, as they often weren’t actual mistakes” and “not quite accurate enough to rely on them as a suggestion”. Some translators also stated that missed errors led them to “disregarding the highlights to focus on checking each sentence”. Despite their high quality, only one editor working with Oracle highlights found highlights helpful in “making the editing process faster and somehow easier”. Taken together, these comments convincingly point to a negative perception of the quality and usefulness of highlights, suggesting that improvement in QE accuracy may not be sufficient to improve QE usefulness in editors’ eyes.
| Question | Italian | Dutch |
| MT outputs were generally of high quality. | ||
| Provided texts were challenging to translate. | ||
| Highlights … | ||
| … were generally accurate in detecting potential issues. | ||
| … were generally useful during editing. | ||
| … improved my editing productivity. | ||
| … improved the quality of my translations. | ||
| … required additional editing effort on my part. | ||
| … influenced my editing choices. | ||
| … helped identify errors I’d have otherwise missed. |
5 Conclusion
This study evaluated the impact of various error-span highlighting modalities, including automatic and human-made ones, on the productivity and quality of human post-editing in a realistic professional setting. Our findings highlight the importance of domain, language and editors’ speed in determining highlights’ effect on productivity and quality, underscoring the need for broad evaluations encompassing diverse settings. The limited gains of human-made highlights over automatic QE and their indistinguishable perception from editors’ assessment indicate that further gains in the accuracy of these techniques might not be the determining factor in improving their integration into post-editing workflows. In particular, future work might explore other directions to further assess and improve the usability of word-level QE highlights, for example, studying their impact on non-professional translators and language learners or combining them with edit suggestions to justify the presence of error spans.
6 Limitations
Our study presents certain limitations that warrant consideration when interpreting its findings and for guiding future research.
Firstly, while we included two domains and translation directions to improve the generalizability of our findings, our results suggest that language and domain play an important role in defining the effectiveness of word-level QE for human post-editing. While we observed mild gains from word-level QE on our tested mid-resourced translation directions (EnglishItalian and EnglishDutch), we expect limited, if any, benefit of such approaches in low-resource languages and domains for which MT systems and QE methods are likely to underperform (Sarti et al., 2022; Zouhar et al., 2024). Furthermore, the domains tested in our study (biomedical and social media posts) provided concrete challenges in the form of specialized terminology and idiomatic expressions, respectively, which are known to hinder the quality of MT outputs (Neves et al., 2024; Bawden and Sagot, 2023). While future work should ensure our findings can be extended to other domains and languages, the limited benefits brought by the tested word-level QE methods in challenging settings suggest a limited usefulness for higher-resource languages and more standard domains such as news or Wiki texts.
Secondly, we acknowledge that several design choices in our evaluation setup, rather than pertaining to the QE methods themselves, may have influenced our results. These include, for instance, the specific procedure for discretizing continuous scores from the Unsupervised method into error spans, and the method of obtaining oracle highlights via majority voting among post-editors. While we believe these choices are justified within the context of our study, their impact on the outcomes cannot be entirely discounted. Future studies might benefit from a more fine-grained assessment of how such low-level decisions influence the perceived accuracy and usability of word-level QE.
Finally, subjective factors such as the translators’ inherent propensity to edit, their prior opinions on the role of MT in post-editing, and their individual editing styles inevitably influenced both quantitative and qualitative assessments in this study. Although we attempted to mitigate these effects by ensuring a controlled evaluation setup for all professional translators and by using averaged judgments for translators working on the same highlight modality, we acknowledge that subjectivity might limit the reproducibility of our findings.
7 Broader Impact and Ethical Considerations
Our study explicitly centers the experience of professional translators, responding to recent calls for user-centered evaluation of translation technologies. By prioritizing translators’ perspectives and productivity, we aim to contribute to methods that complement rather than replace human expertise. Our findings highlight a gap between user perception and measured quality improvements, suggesting that future efforts should focus primarily on improving the usability of these methods in editing interfaces. In particular, new assistive approaches for post-editing should not only strive to increase productivity but rather reduce the cognitive burden associated with post-editing work. This insight is crucial for designing more user-centered quality estimation tools that genuinely support human work. Ultimately, our results suggest that subjective norms across different domains and cultures play an important role in determining the effectiveness of proposed methodologies, underscoring the importance of accounting for human factors when designing such evaluations. All participants in this study were professional translators who provided informed consent. The research protocol ensured anonymity and voluntary participation, with translators recruited and remunerated through professional translation providers. The released materials further promote transparency, enabling other researchers to reproduce and build upon our findings.
Acknowledgements
Authors acknowledge the support of Imminent, which funded the data collection through a Language Technology Grant. Authors are grateful to Translated S.r.l. and Global Textware B.v. and to all translators involved in this study for agreeing to participate in the data collection process. We thank TACL reviewers and editors for their insightful comments. Gabriele Sarti, Grzegorz Chrupała and Arianna Bisazza also acknowledge the support of the Dutch Research Council (NWO) for the project InDeep (NWA.1292.19.399). Arianna Bisazza is supported by the NWO Talent Program (VI.Vidi.221C.009). Ana Guerberof Arenas is supported by the ERC Grant CoG INCREC 101086819.
Appendix A Filtering Details for QE4PE Data
1.
Documents should contain between 4 and 10 segments, each containing 10-100 words (959 docs). This ensures that all documents are roughly uniform in terms of size and complexity to maintain a steady editing flow (Section˜3.5).
2.
The average segment-level QE score predicted by XCOMET-XXL is between 0.3 and 0.95, with no segment below 0.3 (429 docs). This forces segments to have a decent but still imperfect quality, excluding fully wrong translations.
3.
At least 3 and at most 20 errors spans per document, with no more than 30% of words in the document being highlighted (351 docs). This avoids overwhelming the editor with excessive highlighting, while still ensuring error presence.
The same heuristics were applied to both translation directions, selecting only documents matching our criteria in both cases.
| Remove negation (13-6) | |
| English | No significant differences were found with respect to principal diagnoses […] |
| Dutch | Er werden geen significante verschillen significante verschillen gevonden met betrekking tot de belangrijkste diagnoses […] |
| Title literal translation (16-3) | |
| English | The Last of Us is an easy and canonical example of dad-ification. […] |
| Italian | The Last of Us L’ultimo di noi è un esempio facile e canonico di dad-ification. […] |
| Wrong term (48-5) | |
| English | […], , except for alkaline phosphatase. |
| Italian | […], ad eccezione della fosfatasi alcalina chinasi proteica. |
| Target: Seg. Edit Time, 5s bins from 0 to 600s | ||
| Feature | Coeff. | Significance |
| (Intercept) | 1.67 | *** |
| MT Num. Chars | 2.42 | *** |
| Highlight Ratio % | 1.59 | *** |
| Target Lang.: ITA | -0.34 | *** |
| Text Domain: Social | 0.31 | *** |
| Oracle Highlight | -0.79 | . |
| Sup. Highlight | 0.02 | |
| Unsup. Highlight | -0.07 | |
| MT XCOMET QE Score | 0.01 | *** |
| ITA:Oracle | 0.91 | *** |
| ITA:Sup. | 1.18 | *** |
| ITA:Unsup. | 0.48 | *** |
| Social:Oracle | -0.19 | ** |
| Social:Sup. | -0.34 | *** |
| Social:Unsup. | -0.22 | *** |
| Highlight Ratio:Oracle | -0.83 | * |
| Highlight Ratio:Sup. | -1.33 | *** |
| Edit Order | Random Factors | |
| Translator ID | ||
| Segment ID | ||
| Target: % of edited characters in a segment (0-100). | ||
| Feature | Coeff. | Significance |
| (Intercept) | 21.0 | *** |
| MT Num. Chars | 10.3 | *** |
| Highlight Ratio % | 7.1 | *** |
| Target Lang.: ITA | -9.9 | *** |
| Text Domain: Social | 10.9 | *** |
| Oracle Highlight | -5.2 | |
| Sup. Highlight | -4.7 | |
| Unsup. Highlight | -0.9 | |
| ITA:Oracle | 12.2 | *** |
| ITA:Sup. | 15.9 | *** |
| ITA:Unsup. | 13.4 | *** |
| Social:Oracle | 3.5 | *** |
| Social:Sup. | -0.4 | |
| Social:Unsup. | 2.1 | ** |
| Highlight Ratio:Oracle | -0.18 | |
| Highlight Ratio:Sup. | -1.78 | *** |
| Edit Order | Random Factors | |
| Translator ID | ||
| Segment ID | ||
| MT Num. Chars | Zero-Inflation Factors | |
| Target Lang | ||
| Text Domain | ||
| Translator ID | ||
| Identifier | Job | Eng. Lvl | Trans. YoE | Post-edit YoE | Post-edit % | Adv. CAT YoE | MT good/bad for: | Post-edit comment |
| eng-ita-nohigh-fast | Freelance (FT) | C1 | 2-5 | 2-5 | 100% | Often | Good: Productivity, quality, repetitive work. | PE better than from scratch when consistency is needed. |
| eng-ita-nohigh-avg | Freelance (PT) | C1 | >10 | <2 | 20% | Often | Good: Productivity, repetitive work. Bad: less creative. | PE produces unnatural sentences. |
| eng-ita-nohigh-slow | Freelance (PT) | C2 | >10 | 2-5 | 40% | Sometimes | Good: creativity. | Good for time saving. |
| eng-ita-oracle-fast | Freelance (FT) | C2 | 5-10 | 2-5 | 60% | Sometimes | Good: Productivity, repetitive work. Bad: less creative. | Good for productivity, humans always needed. |
| eng-ita-oracle-avg | Freelance (FT) | C2 | 5-10 | 5-10 | 20% | Always | Good: productivity, terminology. | Good for tech docs, not for articulated texts. |
| eng-ita-oracle-slow | Freelance (FT) | C2 | 2-5 | 5-10 | 80% | Always | Good: Productivity, repetitive work. | Useful for consistency and productivity, unless creativity is needed. |
| eng-ita-unsup-fast | Freelance (FT) | C1 | <2 | <2 | 60% | Often | Good: Productivity, terminology. Bad: less creative. | Humans will always be needed in translation. |
| eng-ita-unsup-avg | Freelance (FT) | C1 | >10 | 2-5 | 60% | Often | Good: Productivity, repetitive work. Bad: less creative. | An opportunity for translators. |
| eng-ita-unsup-slow | Freelance (FT) | C1 | 5-10 | 5-10 | 80% | Always | Good: Productivity, repetitive work. Bad: less creative. | Good for focusing on detailed/cultural/creative aspects of translations. |
| eng-ita-sup-fast | Freelance (PT) | C1 | >10 | 2-5 | 40% | Often | Good: Productivity, quality, repetitive work, terminology. | Improves quality and consistency. |
| eng-ita-sup-avg | Freelance (FT) | C1 | >10 | 5-10 | 100% | Always | Good: Productivity, repetitive work. Bad: less creative. | Consistency improved, but less variance means less creativity. |
| eng-ita-sup-slow | Freelance (FT) | C1 | >10 | 2-5 | 20% | Always | Good: Productivity, creativity, quality, repetitive work. | Good for productivity, but does not work on creative texts. |
| eng-nld-nohigh-fast | Freelance (FT) | C1 | >10 | >10 | 40% | Often | Good: Productivity, terminology. Bad: creativity. | Widespread but still too literal |
| eng-nld-nohigh-avg | Freelance (FT) | C2 | >10 | 2-5 | 40% | Always | Good: Repetitive work. Bad: creativity, often wrong, worse quality. | Increase in productivity to save on costs brings down quality. |
| eng-nld-nohigh-slow | Freelance (FT) | C2 | >10 | 5-10 | 100% | Often | Good: Creativity, quality, repetitive work, terminology. | Working with MT can be creative beyond PE. |
| eng-nld-oracle-fast | Freelance (FT) | C1 | 5-10 | 5-10 | 80% | Always | Good: Productivity, quality, repetitive work, terminology. | Good for tech docs and repetition. |
| eng-nld-oracle-avg | Freelance (FT) | C2 | >10 | 2-5 | 40% | Always | Bad: less creative, less productive, often wrong | Bad MT is worse than no MT for specialized domains. |
| eng-nld-oracle-slow | Freelance (FT) | C2 | >10 | 2-5 | 60% | Often | Good: Productivity, repetitive work. Bad: cultural references. | More productivity at the cost of idioms and cultural factors. |
| eng-nld-unsup-fast | Freelance (FT) | C2 | 5-10 | 2-5 | 40% | Often | Good: all. Bad: often wrong, worse quality. | PE makes you less in touch with the texts and often poorly paid. |
| eng-nld-unsup-avg | Freelance (FT) | C2 | 5-10 | 2-5 | 60% | Sometimes | Good: Productivity, quality, repetitive work, terminology. Bad: wrong. | Practical but less effective for longer passages. |
| eng-nld-unsup-slow | Freelance (FT) | C2 | >10 | 2-5 | 40% | Always | Good: repetitive work, productivity, terminology | Improves consistency and productivity if applied well. |
| eng-nld-sup-fast | Freelance (FT) | C2 | >10 | 5-10 | 60% | Often | Good: repetitive work, creativity, terminology | Useful, but worries about job loss |
| eng-nld-sup-avg | Freelance (FT) | C2 | >10 | 10 | 60% | Sometimes | Good: terminology, creativity | Useful for inspiration on better translations |
| eng-nld-sup-slow | Freelance (FT) | C1 | 5-10 | 5-10 | 80% | Always | Good: repetitive work, productivity | Better productivity at the cost of creativity. |
| Identifier | Freq. Issues | MT quality | MT fluency | MT accuracy | High. accurate | High. useful | Interface clear | Task difficult | Highlights statements | |||||
| Speed? | Quality? | Effort? | Influence | Spot errors | Enjoy? | |||||||||
| eng-ita-nohigh-fast | inflection,additions,omissions | 4 | 0.8 | 0.8 | - | - | 5 | 1 | - | - | - | - | - | - |
| eng-ita-nohigh-avg | multiple | 3 | 0.6 | 0.4 | - | - | 2 | 4 | - | - | - | - | - | - |
| eng-ita-nohigh-slow | terminology,omissions | 3 | 0.8 | 0.8 | - | - | 1 | 5 | - | - | - | - | - | - |
| eng-ita-oracle-fast | inflection,terminology | 5 | 0.4 | 0.8 | 4 | 4 | 4 | 5 | 5 | 2 | 1 | 1 | 1 | 4 |
| eng-ita-oracle-avg | syntax,terminology,omissions,no context | 3 | 0.4 | 0.6 | 2 | 1 | 2 | 3 | 1 | 1 | 4 | 1 | 1 | 1 |
| eng-ita-oracle-slow | syntax,no context | 3 | 0.6 | 0.6 | 2 | 2 | 2 | 5 | 1 | 1 | 1 | 1 | 4 | 1 |
| eng-ita-unsup-fast | omissions | 3 | 0.8 | 0.6 | 3 | 2 | 4 | 5 | 3 | 3 | 3 | 2 | 2 | 2 |
| eng-ita-unsup-avg | syntax,terminology,no context | 3 | 0.6 | 0.6 | 3 | 3 | 3 | 5 | 2 | 3 | 2 | 1 | 1 | 3 |
| eng-ita-unsup-slow | syntax,inflection,terminology,omissions | 3 | 0.4 | 0.6 | 2 | 2 | 3 | 4 | 2 | 2 | 3 | 3 | 4 | 4 |
| eng-ita-sup-fast | syntax,terminology,no context | 3 | 0.4 | 0.4 | 2 | 1 | 2 | 2 | 1 | 1 | 3 | 1 | 2 | 2 |
| eng-ita-sup-avg | syntax,terminology,no context | 3 | 0.4 | 0.4 | 2 | 2 | 3 | 5 | 3 | 2 | 4 | 3 | 3 | 4 |
| eng-ita-sup-slow | syntax,terminology,omissions,no context | 3 | 0.6 | 0.6 | 2 | 2 | 1 | 2 | 2 | 1 | 1 | 4 | 4 | 1 |
| eng-nld-nohigh-fast | syntax,terminology,omissions,no context | 3 | 0.2 | 0.4 | - | - | 4 | 4 | - | - | - | - | - | - |
| eng-nld-nohigh-avg | syntax,terminology,omissions,no context | 2 | 0.4 | 0.6 | - | - | 4 | 5 | - | - | - | - | - | - |
| eng-nld-nohigh-slow | terminology,omissions,no context | 2 | 0.2 | 0.4 | - | - | 3 | 5 | - | - | - | - | - | - |
| eng-nld-oracle-fast | syntax,inflection,terminology | 3 | 0.6 | 0.6 | 2 | 1 | 3 | 2 | 2 | 2 | 2 | 1 | 1 | 1 |
| eng-nld-oracle-avg | syntax | 3 | 0.8 | 0.6 | 4 | 3 | 3 | 4 | 3 | 3 | 3 | 3 | 2 | 3 |
| eng-nld-oracle-slow | syntax,terminology | 3 | 0.6 | 0.4 | 3 | 1 | 3 | 4 | 1 | 1 | 1 | 1 | 1 | 3 |
| eng-nld-unsup-fast | terminology,additions,omissions | 3 | 0.6 | 0.8 | 3 | 2 | 4 | 4 | 1 | 3 | 1 | 1 | 2 | 1 |
| eng-nld-unsup-avg | multiple | 3 | 0.6 | 0.6 | 4 | 3 | 2 | 4 | 3 | 3 | 4 | 3 | 2 | 3 |
| eng-nld-unsup-slow | syntax,terminology,omissions | 1 | 0.4 | 0.4 | 2 | 4 | 1 | 4 | 4 | 4 | 3 | 2 | 2 | 3 |
| eng-nld-sup-fast | terminology,omissions,no context | 3 | 0.6 | 0.4 | 2 | 2 | 3 | 5 | 1 | 1 | 5 | 3 | 1 | 1 |
| eng-nld-sup-avg | syntax,additions,no context | 3 | 0.4 | 0.6 | 2 | 2 | 2 | 4 | 1 | 1 | 1 | 1 | 2 | 3 |
| eng-nld-sup-slow | multiple | 5 | 0.8 | 1 | 4 | 3 | 2 | 5 | 3 | 3 | 2 | 2 | 2 | 4 |
| Modalities | EnglishItalian | EnglishDutch | Both | |||||||
| Bio | Social | Both | Bio | Social | Both | Bio | Social | Both | ||
| Oracle and | Sup. | 0.17 | 0.32 | 0.25 | 0.38 | 0.29 | 0.34 | 0.26 | 0.29 | 0.29 |
| Unsup. | 0.14 | 0.30 | 0.20 | 0.31 | 0.27 | 0.28 | 0.22 | 0.29 | 0.24 | |
| Supervised and | Oracle | 0.19 | 0.31 | 0.26 | 0.30 | 0.26 | 0.29 | 0.24 | 0.29 | 0.28 |
| Unsup. | 0.19 | 0.33 | 0.25 | 0.28 | 0.24 | 0.25 | 0.24 | 0.29 | 0.25 | |
| Unsupervised and | Oracle | 0.22 | 0.32 | 0.27 | 0.35 | 0.30 | 0.33 | 0.28 | 0.31 | 0.30 |
| Sup. | 0.22 | 0.37 | 0.30 | 0.39 | 0.27 | 0.33 | 0.30 | 0.31 | 0.32 | |
| # Doc.-Seg. | Error Type | Has Highlight | % Post-edited | |||||
| Oracle | Unsup. | Sup. | No High. | Oracle | Unsup. | Sup. | ||
| 1-8 | Wrong number | nld | Both | Both | 67 | 83 | 83 | 83 |
| 13-6 | Remove negation | ita | Both | Both | 50 | 33 | 33 | 50 |
| 16-3 | Title literal translation | Both | Both | Both | 83 | 100 | 100 | 100 |
| 20-1 | Wrong acronym | nld | Both | ita | 0 | 33 | 33 | 33 |
| 20-7 | Wrong acronym (1) | Neither | Both | Neither | 0 | 58 | 50 | 25 |
| 20-7 | Wrong acronym (2) | nld | Both | ita | 0 | 58 | 50 | 25 |
| 22-1 | Name literal translation | Both | Both | Both | 50 | 50 | 83 | 67 |
| 23-4 | Addition | nld | Both | Neither | 100 | 100 | 83 | 50 |
| 31-2 | Wrong acronym | nld | Both | Neither | 17 | 33 | 17 | 33 |
| 34-7 | Numbers swapped | nld | Both | nld | 17 | 50 | 33 | 67 |
| 37-4 | Verb polarity inverted | Both | Both | Both | 67 | 83 | 67 | 83 |
| 43-5 | Wrong name | Both | Both | Both | 50 | 83 | 67 | 83 |
| 48-5 | Wrong term | nld | Both | nld | 67 | 50 | 83 | 83 |
| Total | 65 | 100 | 62 | 44 | 63 | 60 | 60 | |
| Domain | Modality | |||||||||
| EnglishItalian | ||||||||||
| Biomed. | Random | .12 | - | - /.02 | - /.02 | - /1.0 | - /.11 | - /.13 | - /0.8 | - /.03 |
| No High. | - | .02 | - | - | - | - | - | - | - | |
| Oracle | .08 | .07 | .26/.08 | .05/.02 | 5.2/4.0 | .30/.26 | .06/.08 | 5.0/3.2 | .28/.12 | |
| Unsup. | .16 | .10 | .18/.06 | .08/.02 | 2.2/3.0 | .29/.36 | .14/.15 | 2.0/2.4 | .22/.10 | |
| Sup. | .11 | .12 | .18/.05 | .11/.02 | 1.6/2.5 | .16/.23 | .10/.10 | 1.6/2.3 | .17/.08 | |
| Social | Random | .20 | - | - /.09 | - /.09 | - /1.0 | - /.21 | - /.20 | - /1.0 | - /.13 |
| No High. | - | .09 | - | - | - | - | - | - | - | |
| Oracle | .25 | .20 | .42/.23 | .13/.04 | 3.2/5.7 | .52/.66 | .18/.21 | 2.8/3.1 | .46/.34 | |
| Unsup. | .17 | .18 | .35/.19 | .14/.07 | 2.5/2.7 | .33/.37 | .14/.15 | 2.3/2.4 | .34/.25 | |
| Sup. | .15 | .21 | .38/.23 | .18/.06 | 2.1/3.8 | .27/.39 | .11/.12 | 2.4/3.2 | .32/.29 | |
| EnglishDutch | ||||||||||
| Biomed. | Random | .17 | - | - /.12 | - /.10 | - /1.2 | - /.19 | - /.17 | - /1.1 | - /.15 |
| No High. | - | .10 | - | - | - | - | - | - | - | |
| Oracle | .21 | .08 | .21/.20 | .05/.08 | 4.2/2.5 | .52/.41 | .18/.18 | 2.8/2.2 | .30/.27 | |
| Unsup. | .23 | .09 | .17/.17 | .07/.08 | 2.4/2.1 | .43/.38 | .21/.21 | 2.0/1.8 | .24/.23 | |
| Sup. | .12 | .08 | .20/.21 | .06/.09 | 3.3/2.3 | .30/.25 | .11/.11 | 2.7/2.2 | .24/.23 | |
| Social | Random | .16 | - | - /.22 | - /.19 | - /1.1 | - /.19 | - /.16 | - /1.1 | - /.17 |
| No High. | - | .19 | - | - | - | - | - | - | - | |
| Oracle | .19 | .12 | .33/.39 | .07/.15 | 4.7/2.6 | .54/.39 | .15/.15 | 3.6/2.6 | .41/.39 | |
| Unsup. | .15 | .13 | .25/.33 | .11/.17 | 2.2/1.9 | .30/.26 | .13/.12 | 2.3/2.1 | .27/.29 | |
| Sup. | .12 | .10 | .30/.36 | .08/.17 | 3.7/2.1 | .36/.23 | .10/.10 | 3.6/2.3 | .33/.28 | |
| Domain | Speed | |||||||||
| EnglishItalian | ||||||||||
| Biomed. | Fast | .09 | .04/.01 | .12/.02 | .03/.01 | 4.0/2.0 | .30/.27 | .08/.11 | 3.7/2.4 | .17/.04 |
| Avg. | .10/.05 | .27/.12 | .09/.04 | 3.0/3.0 | .22/.30 | .07/.11 | 3.1/2.7 | .24/.17 | ||
| Slow | .09/.02 | .21/.04 | .08/.01 | 2.6/4.0 | .19/.26 | .07/.11 | 2.7/2.3 | .20/.07 | ||
| Social | Fast | .14 | .11/.07 | .30/.20 | .07/.04 | 4.2/5.0 | .40/.52 | .11/.16 | 3.6/3.2 | .34/.29 |
| Avg. | .23/.14 | .48/.32 | .18/.10 | 2.6/3.2 | .30/.42 | .09/.15 | 3.3/2.8 | .37/.36 | ||
| Slow | .17/.05 | .39/.14 | .14/.03 | 2.7/4.6 | .31/.54 | .11/.17 | 2.8/3.1 | .35/.22 | ||
| EnglishDutch | ||||||||||
| Biomed. | Fast | .14 | .03/.02 | .11/.05 | .02/.01 | 5.5/5.0 | .48/.61 | .13/.18 | 3.6/3.3 | .18/.09 |
| Avg. | .11/.19 | .20/.30 | .10/.17 | 2.0/1.7 | .25/.29 | .13/.16 | 1.9/1.8 | .22/.29 | ||
| Slow | .12/.10 | .26/.23 | .10/.07 | 2.6/3.2 | .29/.42 | .12/.16 | 2.4/2.6 | .27/.30 | ||
| Social | Fast | .12 | .06/.07 | .19/.21 | .04/.04 | 4.7/5.2 | .37/.47 | .10/.13 | 3.7/3.6 | .25/.29 |
| Avg. | .17/.32 | .32/.48 | .15/.29 | 2.1/1.6 | .22/.23 | .10/.12 | 2.2/1.9 | .26/.31 | ||
| Slow | .18/.18 | .38/.40 | .15/.14 | 2.5/2.8 | .25/.34 | .09/.11 | 2.7/3.0 | .30/.37 | ||
| Language | MQM Category | MT | No Highlight | Oracle | Unsupervised | Supervised | |||||
| Maj. | Min. | Maj. | Min. | Maj. | Min. | Maj. | Min. | Maj. | Min. | ||
| Italian | Accuracy - Addition | 0 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 1 |
| Accuracy - Mistranslation | 21 | 22 | 10 | 12 | 4 | 8 | 24 | 17 | 17 | 17 | |
| Accuracy - Inconsistency | 2 | 4 | 1 | 3 | 2 | 2 | 1 | 3 | 0 | 2 | |
| Accuracy - Omission | 2 | 0 | 0 | 0 | 0 | 1 | 4 | 1 | 1 | 2 | |
| Accuracy - Untranslated | 1 | 4 | 1 | 2 | 0 | 1 | 1 | 1 | 3 | 2 | |
| Style - Inconsistent Style | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |
| Style - Readability | 17 | 25 | 5 | 30 | 0 | 12 | 4 | 34 | 1 | 29 | |
| Style - Wrong Register | 0 | 8 | 0 | 3 | 0 | 3 | 1 | 1 | 3 | 2 | |
| Linguistic - Grammar | 6 | 15 | 2 | 16 | 0 | 5 | 3 | 12 | 2 | 12 | |
| Linguistic - Punctuation | 1 | 13 | 0 | 9 | 0 | 3 | 1 | 6 | 0 | 3 | |
| Linguistic - Spelling | 5 | 3 | 0 | 4 | 0 | 3 | 3 | 2 | 0 | 1 | |
| Total | 55 | 95 | 19 | 79 | 6 | 38 | 42 | 77 | 28 | 71 | |
| Dutch | Accuracy - Addition | 0 | 1 | 0 | 2 | 0 | 3 | 0 | 2 | 0 | 1 |
| Accuracy - Mistranslation | 25 | 34 | 18 | 25 | 23 | 27 | 12 | 31 | 16 | 29 | |
| Accuracy - Inconsistency | 0 | 0 | 0 | 2 | 0 | 2 | 0 | 2 | 0 | 5 | |
| Accuracy - Omission | 3 | 1 | 1 | 1 | 2 | 1 | 1 | 1 | 4 | 2 | |
| Accuracy - Untranslated | 4 | 4 | 1 | 1 | 1 | 4 | 1 | 3 | 0 | 2 | |
| Style - Inconsistent Style | 2 | 0 | 0 | 5 | 1 | 7 | 0 | 2 | 0 | 9 | |
| Style - Readability | 1 | 27 | 1 | 20 | 0 | 13 | 2 | 15 | 6 | 41 | |
| Style - Wrong Register | 0 | 2 | 0 | 3 | 0 | 3 | 0 | 1 | 1 | 0 | |
| Linguistic - Grammar | 3 | 19 | 2 | 14 | 3 | 23 | 2 | 6 | 3 | 12 | |
| Linguistic - Punctuation | 0 | 6 | 0 | 3 | 0 | 4 | 0 | 2 | 0 | 3 | |
| Linguistic - Spelling | 1 | 1 | 1 | 1 | 2 | 1 | 0 | 1 | 0 | 0 | |
| Total | 39 | 95 | 24 | 77 | 32 | 88 | 18 | 66 | 30 | 104 | |
| Method | DivEMT | QE4PE | ||||||
| EnIt | EnNl | EnIt | EnNl | |||||
| AP | AU | AP | AU | AP | AU | AP | AU | |
| Logprobs (Fomicheva et al., 2020) | 0.18 | 0.18 | 0.19 | 0.19 | 0.10 | 0.09 | 0.09 | 0.09 |
| Logprobs (Fomicheva et al., 2020, Unsup.) | 0.41 | 0.41 | 0.42 | 0.42 | 0.23 | 0.23 | 0.31 | 0.31 |
| XCOMET-XXL (Guerreiro et al., 2024, Sup.) | 0.16 | 0.23 | 0.19 | 0.28 | ||||
| avg. Oracle single translator | - | - | - | - | 0.53 | 0.73 | 0.55 | 0.75 |
Pre-task
Main task
Post-task
Pre-task
Main task
Post-task